### v1

# An All-digital 65-nm Tsetlin Machine Image Classification Accelerator with 8.6 nJ per MNIST Frame at 60.3k Frames per Second

Abstract
We present an all-digital programmable machine learning accelerator chip for image classification, underpinning on the Tsetlin machine (TM) principles. The TM is a machine learning algorithm founded on propositional logic, utilizing sub-pattern recognition expressions called clauses. The accelerator implements the coalesced TM version with convolution, and classifies booleanized images of 28$\times$28 pixels with 10 categories. A configuration with 128 clauses is used in a highly parallel architecture. Fast clause evaluation is obtained by keeping all clause weights and Tsetlin automata (TA) action signals in registers. The chip is implemented in a 65 nm low-leakage CMOS technology, and occupies an active area of 2.7 mm2. At a clock frequency of 27.8 MHz, the accelerator achieves 60.3k classifications per second, and consumes 8.6 nJ per classification. The latency for classifying a single image is 25.4 $\mu$s which includes system timing overhead. The accelerator achieves 97.42%, 84.54% and 82.55% test accuracies for the datasets MNIST, Fashion-MNIST and Kuzushiji-MNIST, respectively, matching the TM software models.
Index Terms:
Tsetlin machine, Accelerator, Machine learning, Image classification.
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I Introduction
Hardware (HW) accelerators are commonly integrated in modern electronic systems to relieve the burden on system processors from extensive application-specific workloads. Performance and energy efficiency within a smaller form factor are key considerations for accelerator designs, particularly for Internet-of-Things (IoT), where the compute and memory resources of edge nodes typically are limited. In addition, such nodes are often battery-operated, and it is critical to minimize the overall energy footprints for longer battery operational lifetime.
Image classification is a key machine learning (ML) application, with usage scenarios across a wide range of IoT applications [1], such as industrial machine-vision, agricultural monitoring, unmanned aerial vehicles, room occupancy detection for building control systems, attention detection and gesture recognition.
Current methods of accelerating image classification in embedded systems, are predominantly based on convolutional neural networks (CNNs) [2, 3]. CNNs comprise many layers of multiply-and-accumulate (MAC) operations, each with their associated memory resources. As such, they require large number of micro-architectural operations for each classification and have significant energy footprints [3]. The complexity and energy consumption can be reduced by compromising the data precision in CNNs [4]. Binary and ternary activations and weights have been applied in several integrated circuit (IC) solutions that have shown good test accuracy with very low power consumption [5, 6, 7, 8, 9].
Vision Transformer (ViT) [10] solutions obtain the highest score on image classification benchmarks, but they are computationally more intensive than CNNs [11]. ViT-based neural networks utilize a global receptive field that captures correlations between patches in an image [10, 11]. This is in contrast to CNNs, where convolutional layers extract features of an image in a hierarchical manner, with a limited local receptive field. Lightweight Transformer models have been developed, e.g., MobileViT [12], that have been applied for HW implementation [13].
Unlike the numerical models based on neural networks mentioned above, the Tsetlin machine [14] is an ML algorithm founded on propositional logic. During training, parallel conjunctive clauses are defined by independent learning automata. Each clause represents a logical sub-pattern and will output a logical 1 if it matches an input sample, and 0 otherwise. The TM features a single-layer structure with highly interpretable outputs [15, 16, 17], making it well-suited for application areas such as biomedical, cybernetics, IoT, and law, where interpretability is essential. Furthermore, because its primary operations are Boolean, TM solutions are both low-complexity and hardware-efficient. The generation of a TM’s output relies on summing and comparing class sums, with minimal or no need for energy-intensive multiplication operations, even during training. This makes TMs ideal for low-energy application-specific integrated circuit (ASIC) implementations [18]. Additionally, due to their small resource footprint, TMs facilitate hardware solutions that support on-device training and continuous learning [19, 20].
When compared against deep neural networks (DNNs) and other ML algorithms, the TM has demonstrated competitive performance. It has been tested on tabular data [14], natural language [17, 15], images [21, 22, 23] and regression [24]. Through the works in [25, 26, 27] the convergence of the TM’s basic operators has been verified.
The original TM version [14] is called the vanilla TM. For image classification, improved test accuracy is achieved by the convolutional TM (CTM) [21]. The CTM has obtained a peak test accuracy of 99.4% on the MNIST dataset [28]. For the Fashion-MNIST (FMNIST) [29] and Kuzushiji-MNIST (KMNIST) [30] datasets, the peak test accuracies achieved are 91.5% and 96.31% respectively. Ensemble TM approaches have shown promising accuracies for more complex image datasets, such as CIFAR-10 and CIFAR-100, which are currently being investigated. In the case of CIFAR-10, the approach denoted TM Composites [31] has demonstrated a test accuracy of 82.8% [32].
In this paper, we present an all-digital inference accelerator ASIC based on the coalesced TM (CoTM) [22]. The CoTM utilizes a single clause pool, in contrast to the vanilla TM which uses one pool of clauses per class. As we operate the CoTM with convolution, we here denote it as a convolutional CoTM (ConvCoTM). For configurations with few clauses, the ConvCoTM achieves better accuracy than the vanilla CTM, which makes the ConvCoTM especially attractive for HW solutions with constrained resources.
The main motivation behind the development of this ASIC was to demonstrate the HW friendliness and power efficiency of a ConvCoTM-based accelerator. The main contributions of this paper are:
  • The accelerator is the first reported manufactured ASIC based on the CoTM. It is also the first TM ASIC where convolution is utilized.
  • The ConvCoTM accelerator core operates with an energy per classification (EPC) of 8.6 nJ. To the best of the authors’ knowledge, this is the lowest reported EPC for manufactured ASICs that operate on a similar test dataset and with comparable test accuracy.
  • The accelerator has low latency, essential for real-time operation in low-power systems with strict requirements.
  • The ConvCoTM solution is a fully digital, synchronous design compatible with standard digital workflows.
The remainder of the paper is organized as follows: Section II summarizes related work. Section III covers TM and CoTM background, including convolution operation. Section IV details the ASIC’s architecture and building blocks. Measurement results are presented in Section V, followed by a discussion in Section VI before we conclude the paper.
II Related Work
In this section, we describe state-of-the-art works on low-power inference accelerators for image classification, as well as some reported TM-based HW solutions.
For low-power neural networks, digital Binary Neural Networks (BNNs) are widely utilized. A BNN-based accelerator IC module is detailed in  [33]. This module was later included in a complete system-on-chip (SoC) [6]. When operating on the CIFAR-10 [34] dataset, the SoC achieves 15.4 inferences/s and has a peak power envelope of $674~{}\mu$W. A 22 nm fully depleted silicon on insulator (FD-SOI) technology was used for the IC implementation.
In [7] a digital 10 nm FinFET CMOS solution is reported. Utilizing a BNN algorithm, it achieves a peak energy efficiency of 617 TOPS/W (Tera operations per second per watt). It has a power consumption of 5.6 mW and obtains an accuracy of 86% when operating on the CIFAR-10 dataset.
An architecture for a binary-weight CNN is described in [8]. Images of size $224\times 224\times 3$ are classified at a rate of 11 frames per second (FPS). The technology used is 65 nm CMOS and simulations show a power consumption of 895 $\mu$W at 0.6 V.
The accelerator ASIC in [35] is manufactured in 65 nm CMOS. Its classification rates for Imagenet [36] are 35/0.7 FPS for the AlexNet and VGG-16 models respectively. The ASIC consumes 278 mW/236 mW with corresponding EPC values of 7.9 mW/0.4 mJ. The successor to [35] is the work described in [37]. Here, additional steps are taken to reduce energy consumption further. In particular, it reduces the movement of all data types during the various operational steps of the implemented CNN algorithm. A highly flexible on-chip network is utilized and sparse data are processed efficiently. When operating on $224\times 224\times 3$ images from Imagenet, simulations show a power consumption of 1.44 mW at 1470.6 FPS. The technology used is 65 nm CMOS.
In-memory-computing (IMC) solutions can provide very high energy efficiency, as the bottleneck related to memory access is relieved. IMC has especially been utilized for analog/mixed-signal ML accelerator designs. An IMC mixed-signal 28 nm CMOS BNN processor is described in [5]. This accelerator achieves an EPC of 3.8 $\mu$J when operating on CIFAR-10. The classification rate is 237 FPS with a power consumption of 0.9 mW.
In [9] a charge-domain IMC ternary neural network (TNN) classifier is described. The main advantage of the ternary approach is a reduction of the required operations per classification with 3.9$\times$ compared to a BNN model. The accelerator classifies MNIST images at a rate of 549 FPS. It has very low power consumption, only 96 $\mu$W, which corresponds to 0.18 $\mu$J per classification. The test accuracy achieved is 97.1%.
Neuromorphic approaches, utilizing spiking neural networks (SNNs) have achieved very low energy consumption. In [38] a mixed-signal neuromorphic SNN IC is reported, which has been manufactured in a 65 nm CMOS technology. It achieves a test accuracy on MNIST of 95.35%, a power consumption of 0.517 mW and an EPC of 12.92 nJ. Simulation results of a 28 nm CMOS charge-domain computing SNN solution are reported in [39]. The accelerator achieves an EPC of 15.09 nJ, a latency of 0.46 $\mu$s and a test accuracy on MNIST of 94.81%. In [40] simulation results of a clock-free, event-driven, mixed-signal 28 nm IMC SNN are described. On MNIST it achieves a test accuracy of 96.92% and 1.38 nJ EPC.
For the TM-based hardware solutions, the first reported IC is described in [18]. It utilizes the vanilla TM architecture and operates on a 3-class machine learning (ML) problem. The IC is implemented in a 65 nm CMOS technology and supports both training and inference. In [19], the first TM-based HW accelerator with convolution is reported. It is an FPGA solution which operates on a 2-class pattern recognition problem (two-dimensional noisy XOR [21]) in 4$\times$4 Boolean images, and employs a 2$\times$2 convolution window. Full on-device training is implemented.
The FPGA solution reported in [41] is based on the vanilla TM and stores and processes only the include signals of the TA actions, utilizing the typical TM model’s high sparsity. It therefore represents a highly compact implementation when it comes to HW resources. The design operates on the MNIST dataset, with sequential processing of each clause and each literal.
In [20] the first reported ConvCoTM-based HW solution is presented. It is an FPGA solution that classifies $28\times 28$ images with 10 categories, and includes complete on-device training. For the MNIST dataset a test accuracy of 97.6% is achieved, with a classification rate of 134k samples per second, and an EPC of 13.3 $\mu$J.
Several articles describe the design and simulation of HW solutions for TM accelerators. In [42, 43, 44] low-latency asynchronous TM solutions for inference and training are presented. A mixed-signal IMC TM solution is described in [45] with an EPC for MNIST of 13.9nJ. It is an inference solution with a Boolean-to-current architecture, and resistive RAM (ReRAM) transistor cells are employed to eliminate the need for energy-hungry analog-to-digital and digital-to-analog conversions. Another IMC TM concept based on Y-flash cells is reported in [46].
Based on the above literature review, developing a ConvCoTM ASIC solution is a critical step toward achieving higher energy efficiency, performance, and scalability for TM-based hardware accelerators. Existing FPGA implementations, while effective, are limited in energy efficiency. Comparatively, in-memory computing (IMC) and binary neural network (BNN)-based ASICs demonstrate significantly lower energy consumption and higher operational efficiency, underscoring the potential benefits of an ASIC approach. By leveraging a standard digital design flow and CMOS technology, an ASIC solution for ConvCoTM could deliver real-time, power-efficient classification. This would position ConvCoTM as a competitive alternative to CNNs and BNNs in resource-constrained environments.
III TM and CoTM Background
The basic concepts of a vanilla TM [14] are described here, followed by a description of general CoTM inference [22]. Later, we explain how to apply convolution to the CoTM. As the emphasis in this paper is on inference solutions, we do not include details about CoTM/ConvCoTM training, except for a brief description of the Tsetlin automaton (TA), which is the core learning element in a TM. TM training is covered in detail in the original TM, CTM and CoTM references [14, 21, 22]. A compact description of ConvCoTM training can be found in [20].
III-A General TM Concepts
The input feature vector to a TM consists of $o$ propositional Boolean variables, $x_{u}\in\{0,1\}^{o}$, where $u\in\{0,\ldots,o-1\}$. A new input vector named the literals, L, is generated by appending the negated variables to the input, as shown in Eq. (1). Thus, there are in total $2o$ literals.
$L=[l_{0},l_{1},\dots,l_{2o-1}]\\ =[x_{0},\dots,x_{o-1},\neg x_{0},\dots,\neg x_{o-1}]$ (1)
The Tsetlin Automaton (TA) of two-action type [47] constitutes the basic learning element in a TM. The TA will either include or exclude a literal in a conjunctive clause, depending on its state. One team of TAs is applied per clause, and the different TA actions are obtained during the training process which involves feedback mechanisms to each TA [14]. For an inference-only solution, as reported in this paper, we need only the TA action signals from the trained model, not the complete TAs.
In a TM the number of clauses, $n$, is a user specified integer. The output of a single clause, $c_{j}$, where $j\in\{0,\ldots,n-1\}$, is given by
$c_{j}=\bigwedge_{k\in I_{j}}l_{k},$ (2)
where $l_{k}$ is the literal with index k. k belongs to $I_{j}\subseteq\{0,....,2o-1\}$, where $I_{j}$ is the set of literal indexes for clause $j$ where the TA action is include [14].
III-B CoTM
A vanilla TM employs one clause pool per class, and the clauses of each TM are grouped in two: positive and negative. In contrast, a CoTM [22] applies a single clause pool, common for all $m$ outputs/classes. Different sets of weights, $w_{i,j}$, are applied to the clause outputs, $\{c_{0},\dots,c_{n-1}\}$, to generate the class sums:
$v_{i}=\sum_{j=0}^{n-1}w_{i,j}c_{j},$ (3)
where $i\in\{0,\dots,m-1\}$ is the class index.
The clause weights in Eq. (3) obtain their values during training and can have both positive and negative polarity. It should be noted that no multiplications are needed to implement Eq. (3), only additions, as a clause can output only 0 or 1. The predicted class, $\hat{y}$, is determined by the argmax operator, as given by Eq. (4).
$\hat{y}=\text{argmax}_{i}\{v_{i}\},~{}i\in\{0,\ldots,m-1\}.$ (4)
III-C CoTM and Convolution
One can improve the image classification test accuracy by applying convolution [21] to the CoTM. In addition, the number of features that need to be processed simultaneously, is lowered with convolution, which again reduces the model size and the complexity of a design. However, the benefits of convolution come at the expense of increased processing time, which is due to the need to sequentially process the different patches of the input sample.
For image classification, the input data samples to a ConvCoTM, are a set of images, each of dimensions $X\times Y$, with $Z$ channels. The different patches of the image are generated by applying a sliding window of size $W_{X}\times W_{Y}$, with $Z$ channels. The convolution window will be evaluated $B=B_{X}\times B_{Y}$ times across the image, where $B_{X}=1+(X-W_{X})/d_{x}$, $B_{Y}=1+(Y-W_{Y})/d_{y}$, and $d_{x}$ and $d_{y}$ are the stride values of the convolution window in the $X$ and $Y$ directions respectively [21]. For each window position, the number of Boolean features is given by
$N_{F}=W_{X}\times W_{Y}\times Z\times U+(Y-W_{Y})+(X-W_{X}).$ (5)
In Eq. (5), $U$ is the number of bits used for thermometer encoding [48] of the value of a single pixel (one-hot encoding can also be used). Such encoding is required because a TM operates on Boolean variables, not on binary numbers. The term $(Y-W_{Y})+(X-W_{Y})$ in Eq. (5) represents the number of bits that encode the patch’s position [22, 49] in the $Y$ and $X$ directions respectively. Also here we apply thermometer encoding. As shown in Eq. (1), the ConvCoTM also takes as input the negated versions of the features to form the literals per patch.
III-D Datasets and Booleanization
The widely used MNIST [28] dataset consists of handwritten digits. FMNIST [29] is more challenging than MNIST and includes images from a Zalando catalog, such as t-shirts, sandals, and shoes. KMNIST [30] consists of cursive Japanese characters, and is also more demanding than MNIST. Each of these datasets includes images of $28\times 28$ pixels organized in 60k training samples and 10k test samples. There are 10 classes within each dataset. As the images are of single-channel (greyscale) type, $Z=1$.
The pixel values range from 0 to 255. For MNIST, they are converted into Boolean variables (i.e., booleanized) with simple thresholding, where pixel values larger than 75 are replaced with 1, and 0 otherwise [21]. For FMNIST and KMNIST, the booleanization is performed with an adaptive Gaussian thresholding procedure [21]. Fig. 1 shows one example of an original image and the corresponding booleanized image from each of the three datasets. For the booleanized images the color white represents a logical 1.
Figure 1: Examples of MNIST, FMNIST and KMNIST image booleanization.
For the booleanization, we use one bit per pixel, i.e., $U=1$ for all three datasets. The convolution window applied has dimensions $W_{X}=W_{Y}=10$, and a step size of 1 is used in both directions. The number of feature bits per patch is therefore 100 bits from the convolution window plus 36 bits for the appended patch position information, ref. Eq. (5). The number of literals for each patch is 272, i.e., two times the number of features.
III-E TM and Convolution - Inference
During inference, each clause in a ConvCoTM is evaluated $B$ times per image. After finishing the convolution, a clause will output 1 if it has recognized a pattern at least once in any of the $B$ patches for a given image [21]. The clause output is given by the sequential OR-function in Eq. (6), where $c_{j}$ is the $j$’th clause of the ConvCoTM, and $c_{j}^{b}$ is the $b$’th output of this clause obtained during the window sliding (patch generation).
$c_{j}=\bigvee_{b=0}^{B-1}c_{j}^{b},$ (6)
Class sum generation and class prediction are similar to the case without convolution, and follow Eq. (3) and Eq. (4).
IV Design Overview
The block diagram of the ConvCoTM accelerator is shown in Fig. 2. The accelerator includes all circuitry required for performing inference, and only minimal interaction with the system processor is required. The three main modules are i) a data interface for transfer of model parameters and image samples, ii) a module for model storage and iii) an inference core that performs classification of an image.
Figure 2: ConvCoTM accelerator ASIC block diagram.
The main configuration parameter of a TM solution is the number of clauses. Due to chip area restrictions, we decided to implement a configuration with 128 clauses. The following subsections describe the different sub-modules and the accelerator’s operation in detail.
IV-A Interface to the System Processor
An external system processor is required for i) applying a few single-bit control signals to the accelerator, ii) feeding data to the accelerator and iii) reading interrupt signals and class predictions from the accelerator. All other operations are performed by the accelerator.
The data interface towards the external system processor, is an 8-bit parallel type inspired by the ARM AXI Stream interface [50]. It is used for transferring model data to the accelerator during load model mode, and for transferring images in inference mode. When a classification operation is finished, the accelerator outputs an interrupt. In addition, it outputs a separate 8-bit signal consisting of the predicted class (4 bits) together with the true label of the sample (4 bits).
IV-B ConvCoTM Model Registers
The model of a given CoTM configuration consists of: i) the TA action signals (include/exclude signals) per clause and ii) the signed weights per clause per class. The TA action signals are stored in registers, readily available for simultaneous clause evaluation. These signals require $272\times 128=34816$ D flip-flops (DFFs). Registers are also applied for the clause weights per class. For this ASIC, eight bits are allocated per weight, with two’s complement representation. The number of DFFs for storing the weights is $10\times 128\times 8=10240$, and the complete model size used by the accelerator is 45056 bits, i.e., 5632 bytes.
IV-C Image Data Buffer and Patch Generation
A data buffer is included with room for two complete $28\times 28$ booleanized images. While one image is being processed, another image can be transferred from the external system processor to the other image buffer, thereby improving throughput. We denote this as the continuous classification mode. A complete booleanized image requires $28\times 28/8=98$ bytes. In addition, one byte is allocated for the image label (0 to 9).
The patch generation module consists of a register with 10 rows, each with 28 DFFs. At the start of processing a new sample, the first 10 rows of an image are loaded into the register, and the convolution window starts at the leftmost position, which has coordinates (0,0).
During the convolution, the window is moved to the right with increments of one. As described in subsection III-C, the position of the window is thermometer encoded. There are 19 different $x$-coordinates for the window, requiring 18 bits. When the window has been evaluated at the rightmost $x$-position (0,18), the contents of all 10 rows are shifted upwards with one step, and the next image datarow is loaded into the lowermost row. Then the window starts again at the leftmost position, now with patch coordinates (1,0), and moves to the right in steps of one. This procedure is repeated until all 28 datarows of the image have been processed. Also for the $x$-coordinate, there are 19 different positions. During a complete convolution, $19\times 19=361$ patches are generated.
IV-D Clause Pool
The ConvCoTM is configured with 128 clauses that operate in parallel. Each of the 128 clauses takes the literals, $[l_{0},l_{1},\dots,l_{271}]$, as input, as well as their corresponding TA action signals, $[i_{0},i_{1},\dots,i_{271}]$, from the model register. Fig. 3 shows a simplified circuit diagram of the clause logic. A literal will be included in the logical AND expression, that constitutes the clause, if its corresponding TA action signal is high. A special situation occurs when a clause is empty, i.e., when there are no include TA actions. In this case, additional logic will set the signal Empty high, forcing $c_{j}^{b}$ low.
The DFF in Fig. 3 constitutes a single-bit register for clause $j$. During the convolution, its content, $c_{j}$, is ORed with the combinational clause value, $c_{j}^{b}$, of the current patch, thus implementing the sequential OR function in Eq. (6). Before starting a new convolution (patch generation), the DFF is reset.
The signal $c_{j}$ is also fed back to the input of the OR-gates, as there is no need for the combinational logic to evaluate a clause for more patches if $c_{j}$ already is high. This feedback mechanism, hereinafter denoted the clause switching reduction feedback (CSRF), reduces the digital switching in the combinational part of the clause logic. Simulations on MNIST image classification with CSRF enabled, showed an average of 50% reduction in the toggling rate of $c_{j}^{b}$ for each clause. The accelerator ASIC has a dedicated pin that can enable or disable the CSRF.
Figure 3: Circuitry for a single clause. Literals are denoted with $l$ and TA-actions with $i$.
In this accelerator ASIC we apply a single convolution window. If increased classification rate is wanted, one can use several convolution windows that operate in parallel. In this case, the combinational clause logic that generates the signal $c_{j}^{b}$ in  Fig. 3, would have to be replicated for each convolution window.
IV-E Inference
The inference operation starts with loading an image into the data buffer, followed by a convolution process, as described in Subsection IV-C. When finished, each clause value, $c_{j}$, is weighted, for each class, according to Eq. (3).
All ten class sums are generated in parallel, by adders configured in a tree structure and utilizing a three-stage pipeline. Finally, an argmax module [18] selects the class with the highest class sum, as the predicted class, $\hat{y}$. Fig. 4 shows a simplified state diagram of the accelerator.
Figure 4: Simplified accelerator state diagram.
The accelerator’s latency, measured from when the system processor initiates the first byte transfer of the image, until a prediction result is available, is 471 clock cycles. This includes 99 clock cycles for transferring the 98 image bytes and the label byte, and 372 clock cycles for the patch generation and class prediction.
The accelerator can also operate in continuous mode, where a new sample is loaded into the image buffer while the current sample is being processed, see Subsection IV-C. This increases the throughput, as image samples can be processed every 372’th clock cycle. Any timing overhead in the system processor will add to the total latency.
IV-F Clock Domains
The key to reducing the digital switching power of the accelerator ASIC lies at the top level of the ASIC architecture. Two separate clock domains are utilized, each with its own dedicated clock pin, assigned to i) the model part and ii) the inference core.
When a model has been transferred from the system processor to the ASIC, the TA action signals and the weights are available in registers which do not need further clocking. By stopping the clock to the model part when operating in inference mode, the digital switching power is reduced significantly, as the model part constitutes approximately 90% of the accelerator’s DFFs. To further reduce power consumption, standard clock gating is applied for the inference module. This ensures, for instance, that the pipelined registers associated with the class sum generation are enabled and clocked only for 4 clock cycles per classification cycle. The clock-gating can be enabled/disabled by an external pin.
V Implementation and Measurement Results
The ConvCoTM accelerator was designed in VHDL and first implemented and verified on a Xilinx ZCU104 FPGA development board. The ASIC was then implemented in a 65 nm low-leakage CMOS technology from UMC, and packaged in a JLCC44 ceramic package. Fig. 5 shows a photograph and a layout plot of the ASIC. The accelerator’s core area is 2.7 mm2.
Figure 5: Photo and layout plot of the ConvCoTM accelerator chip.
A test PCB was developed for the ASIC. This board was connected to a Digilent Zybo Z7-20 development board equipped with a Xilinx Zynq XC7Z020 FPGA, as shown in Fig. 6. One of the FPGA’s ARM9 cores was configured as system processor, and the FPGA board provided clocks, model data, image data samples and control signals to the ASIC accelerator. The processor also read back accelerator status/interrupt signals and the predicted/true class per sample. To interface with the FPGA board, the ASIC’s input/output (IO) ports operated at 3.3 V supply voltage. The accelerator core operated from 0.82 V to 1.2 V.
Figure 6: Test setup with the ConvCoTM ASIC test board connected to the FPGA development board.
For each of the datasets MNIST, FMNIST and KMNIST, the Tsetlin Machine Unified (TMU) SW-version [51] of the ConvCoTM was trained to find suitable models. Maximum/minimum limits were set on the clause weights to fit with the allocated 8-bits per weight in the accelerator. For simplicity, the images from the test dataset were preprocessed (booleanized), as described in subsection III-D, and included in the FPGA processor’s test program.
Model data and image samples were applied from the FPGA to the accelerator via the AXI Stream interface. We utilized DMA functionality on the FPGA to enable fast data transfer. The maximum classification rate achieved, including system processor timing overhead, was 60.3k images per second in continuous mode, with an accelerator clock frequency of 27.8MHz. The speed limiting factor was the electrical interface between the FPGA and the ASIC. The latency when classifying a single image was 25.4 $\mu$s, including the transfer time of the image to the accelerator as well as timing overhead in the system processor.
The achieved test accuracies of the datasets MNIST, FMNIST and KMNIST were 97.42%, 84.54% and  82.55% respectively. These results are independent of the accelerator clock frequency and are exactly in accordance with the performance of the models obtained from the SW simulations.
In a real application, the ConvCoTM design would typically be included as a processor peripheral module in an SoC. Therefore, we considered the power consumption of the accelerator core as the most interesting performance power parameter, and our measurements excluded energy consumed by the digital 3.3 V IOs of the ASIC. Table I summarizes the characteristics and performance of the accelerator.
TABLE I: The ConvCoTM accelerator ASIC’s main characteristics and performance
Parameter / Characteristics Value / Description
Technology 65 nm low-leakage CMOS from UMC
Chip area (full die) 3.5 mm2
Chip area (core) 2.7 mm2
Gatecount (core) 201k cells
including 52k DFFs
Power consumption 1.15 mW a,c
(accelerator core only) 0.52 mW a,d
81 $\mu$W b,c
21 $\mu$W b,d
Classification rate 60.3k images/s a
(including system overhead) 2.27k images/s b
Energy per classification 19.1 nJ a,c
(accelerator core only) 8.6 nJ a,d
35.3 nJ b,c
9.6 nJ b,d
Latency (single image classi-, 25.4 $\mu$s a
fication including data transfer) 0.66 ms b
Test accuracy 97.42% (MNIST)
84.54% (FMNIST)
82.55% (KMNIST)
a 27.8 MHz clock frequency. b 1.0 MHz clock frequency. c The accelerator core supply voltage (vdd) is 1.20 V. d vdd=0.82 V.
For power consumption measurements, a Joulescope JS220 precision energy analyzer was used. The power consumption during inference was measured while the accelerator was set in a test mode with repeated classifications of the full 10k sample dataset. The power consumed by the accelerator core with 1.20 V supply voltage and a clock frequency of 27.8 MHz was 1.15 mW. This corresponds to an EPC of 19.1 nJ. With a supply voltage of 0.82 V, the power consumption was 0.52 mW and the corresponding EPC was 8.6 nJ. With 1.2 V core supply voltage and a clock frequency of 1 MHz, the achieved classification rate was 2.27k images/second at a power consumption of 81 $\mu$W.
All power and energy measurement results of the ASIC are average values, and were performed with clock-gating and CSRF enabled (as shown in Fig. 3). When operating at 27.8 MHz, clock-gating reduced the power consumption by approximately 60%, while the CSRF alone provided less than 1% power reduction.
VI Discussions
We consider the ConvCoTM solution with 128 clauses as a small configuration. In light of this, we are satisfied with the obtained test accuracy which is comparable with other low-complexity and low-power solutions. Table II shows the performance and properties of state-of-the-art low-power IC implementations for image classification. Inspired by the TinyML [52] benchmarks for embedded systems we specify the classification rate, the peak power consumption and the EPC. These are system parameters which are independent of the accelerator model and architecture [53]. However, when used for comparing solutions, one must be careful to specify the ML task performed. The ICs shown in Table II operate on images and datasets that range from low to high complexity. Accelerators that operate on CIFAR-10 and other more complex images, typically have lower classification rates and higher EPC compared to accelerators that operate on $28\times 28$ images.
TABLE II: Comparison of the ConvCoTM IC accelerator with previous works.
This work TCAS-II’23 Yejun [38] JSSC’23       Yang [9] TCAS-I’20 Mauro [6] JSSC’21       Knag [7] TCAS-I’20 Bankman [5] JSSC’17 Chen [35] JETCAS’19 Chen [37]
Technology 65 nm CMOS 65 nm CMOS 40 nm CMOS 22 nm FD-SOI 10 nm FinFET 28 nm CMOS 65 nm CMOS 65 nm CMOS
Active area 2.7 mm2 0.57 mm2 0.98 mm2 2.3 mm2 0.39 mm2 4.6 mm2 12.3 mm2 Not stated
Algorithm ConvCoTM SNN Ternary CNN BNN BNN BNN CNN Sparse DNN
Design type Digital Neuromorphic mixed-signal IMC, mixed-signal Digital (SoC) Digital IMC, mixed-signal Digital Digital
Image dataset MNIST, FMNIST, KMNIST MNIST MNIST CIFAR-10 CIFAR-10 CIFAR-10 Imagenet Imagenet
Measured test accuracy 97.42%, 84.54%, 82.55% 95.35% 97.1% 99% of nominal accuracy 86% 86% Not stated 79.6% i           79.7% j
Classifications per second 60.3k a           2.27k b 233k (1.2 V)           40k (0.7 V) 549 15.4 Not stated 237 35 g           0.7 h 278 f,i           1470 f,j
Latency e 25.4 $\mu$s a           0.66 ms b Not stated Not stated Not stated Not stated Not stated 115.3 ms Not stated
Inference power 1.15 mW a,c 0.52 mW a,d 81 $\mu$b,c 21 $\mu$b,d 9.37 mW (1.2 V) 0.517 mW (0.7 V) 96 $\mu$W 674 $\mu$W 5.6 mW (13 MHz) 607 mW (622 MHz) 0.9 mW 278 mW g 236 mW h Not stated
Energy per classification (EPC) 19.1 nJ a,c           8.6 nJ a,d           35.3 nJ b,c           9.6 nJ b,d 40.17 nJ (1.2 V) 12.92 nJ (0.7 V) 0.18 $\mu$J 43.8 $\mu$J Not stated 3.8 $\mu$J 7.9 mJ g           337 mJ h 1.5 mJ f,i 0.4 mJ f,j
a 27.8 MHz clock frequency. b 1.0 MHz clock frequency. c The accelerator core supply voltage (vdd) is 1.20 V. d vdd=0.82 V.e For processing of a single image sample, including data transfer. f Simulated results. g AlexNet. h VGG-16. i Sparse AlexNet. j Mobilenet.
The mixed-signal IMC-based solution in [9] operates with a very low peak power consumption of 96 $\mu$W and consumes 0.18 $\mu$J per classification on MNIST with a test accuracy of  97.1%. Its classification rate is 549. In [38] a mixed-signal SNN IC accelerator achieves a very low EPC of 12.92 nJ with 0.7 V supply voltage. Operating at 1.2 V the EPC is 40.17 nJ. It utilizes a time-domain echo state network (ESN) [54] approach and thus requires non-standard interfacing with the surrounding system’s resources. It obtains a test accuracy on MNIST of 95.35%.
To the authors’ knowledge, the ConvCoTM accelerator core’s EPC of 8.6 nJ at 0.82 V operating voltage is the lowest reported for a manufactured ASIC that classifies $28\times 28$ greyscale images and which has comparable test accuracy. At 1.2 V operating voltage it obtains an EPC of 19.1 nJ, which is also very low. As the ConvCoTM ASIC is fully digital, the classification rate and thereby also the power consumption, can easily be scaled by adjusting the clock frequency. E.g, with a clock frequency of 1 MHz, the power consumption achieved is only 81 $\mu$W. As the accelerator is all-digital, the test accuracy is independent of the clock frequency.
Several analog/mixed-signal IC accelerators obtain very low power consumption [38, 9, 5]. Such solutions are susceptible to variations in process parameters, supply voltage and operating temperature (PVT), which can affect the performance. All-digital solutions avoid such non-ideal effects [6, 7, 8]. In addition, they are easier to integrate in larger SoCs, and can be scaled and re-targeted to other technologies. This applies for CNNs as well as for TMs.
In the ConvCoTM accelerator, the combinational clause logic draws only a small amount of energy compared to the clock tree of the inference-core DFFs. This is the main reason why the CSRF technique, see Subsection IV-D, did not provide a greater reduction in EPC. Other convolutional TM architectures, that operate at higher clock frequencies, process more literals in parallel, and utilize sparsity in TM models, can potentially benefit more from the CSRF technique.
The ConvCoTM accelerator’s architecture described in this paper is highly suited for scaling. It can be modified to operate on larger images including 3-channel (RGB) images, e.g., based on the TM Composites architecture [31, 32]. From the very low energy consumption obtained for our ConvCoTM accelerator, we envisage that a TM Composites-based solution will perform well compared to previous works.
VII Conclusion
In this work, an ASIC implementation of a ConvCoTM accelerator for inference is presented. It operates on booleanized images of 28$\times$28 pixels, and samples are classified into 10 categories. A pool of 128 clauses, i.e., pattern recognition logical expressions, is applied for the accelerator. The solution is fully programmable, and achieves a test accuracy of 97.42% on MNIST. With a clock frequency of 27.8 MHz, the accelerator’s throughput is 60.3k images per second, and with an operating voltage of 0.82 V, the energy consumed by the accelerator core per classification is 8.6 nJ. To the authors’ best knowledge this is lower than for previously reported fabricated ICs for the same classification task and for similar test accuracy.
The ASIC demonstrates the simplicity, hardware-friendliness and power-efficiency of TM solutions. We believe the TM represents an attractive ML solution for low-power edge nodes in IoT systems. Furthermore, in a TM ASIC, full on-device training capabilities can be included, enabling adaptive edge solutions.
Acknowledgments
The authors would like to thank Adrian Wheeldon (Literal Labs, UK) and Jordan Morris (UK) for valuable general ASIC implementation discussions. Thank you also to Joel Trickey (STFC, UK) for support with IC design backend tools, and to Alberto Pagotto (imec.IC-Link, Belgium) for help during the ASIC tapeout phase. Finally, we would like to thank Geir Rune Angell (Westcontrol, Norway) for the design of the ASIC test board.

### v2

@@
-An All-digital 65-nm Tsetlin Machine Image Classification Accelerator with 8.6 nJ per MNIST Frame at 60.3k Frames per Second
+An All-digital 8.6-nJ/Frame 65-nm Tsetlin Machine Image Classification Accelerator
 Abstract
 We present an all-digital programmable machine learning accelerator chip for image classification, underpinning on the Tsetlin machine (TM) principles.
-The TM is a machine learning algorithm founded on propositional logic, utilizing sub-pattern recognition expressions called clauses.
+The TM is an emerging machine learning algorithm founded on propositional logic, utilizing sub-pattern recognition expressions called clauses.
 The accelerator implements the coalesced TM version with convolution, and classifies booleanized images of 28$\times$28 pixels with 10 categories.
 A configuration with 128 clauses is used in a highly parallel architecture.
-Fast clause evaluation is obtained by keeping all clause weights and Tsetlin automata (TA) action signals in registers.
+Fast clause evaluation is achieved by keeping all clause weights and Tsetlin automata (TA) action signals in registers.
 The chip is implemented in a 65 nm low-leakage CMOS technology, and occupies an active area of 2.7 mm2.
 At a clock frequency of 27.8 MHz, the accelerator achieves 60.3k classifications per second, and consumes 8.6 nJ per classification.
+This demonstrates the energy-efficiency of the TM, which was the main motivation for developing this chip.
 The latency for classifying a single image is 25.4 $\mu$s which includes system timing overhead.
@@
 Hardware (HW) accelerators are commonly integrated in modern electronic systems to relieve the burden on system processors from extensive application-specific workloads.
-Performance and energy efficiency within a smaller form factor are key considerations for accelerator designs, particularly for Internet-of-Things (IoT), where the compute and memory resources of edge nodes typically are limited.
+Performance and energy-efficiency within a smaller form factor are key considerations for accelerator designs, particularly for Internet-of-Things (IoT), where the compute and memory resources of edge nodes typically are limited.
 In addition, such nodes are often battery-operated, and it is critical to minimize the overall energy footprints for longer battery operational lifetime.
 Image classification is a key machine learning (ML) application, with usage scenarios across a wide range of IoT applications [1], such as industrial machine-vision, agricultural monitoring, unmanned aerial vehicles, room occupancy detection for building control systems, attention detection and gesture recognition.
-Current methods of accelerating image classification in embedded systems, are predominantly based on convolutional neural networks (CNNs) [2, 3].
+Current methods for accelerating image classification in embedded systems, are predominantly based on convolutional neural networks (CNNs) [2, 3].
 CNNs comprise many layers of multiply-and-accumulate (MAC) operations, each with their associated memory resources.
 As such, they require large number of micro-architectural operations for each classification and have significant energy footprints [3].
-The complexity and energy consumption can be reduced by compromising the data precision in CNNs [4].
+The complexity and energy consumption can be reduced by compromising the data precision [4].
 Binary and ternary activations and weights have been applied in several integrated circuit (IC) solutions that have shown good test accuracy with very low power consumption [5, 6, 7, 8, 9].
-Vision Transformer (ViT) [10] solutions obtain the highest score on image classification benchmarks, but they are computationally more intensive than CNNs [11].
-ViT-based neural networks utilize a global receptive field that captures correlations between patches in an image [10, 11].
-This is in contrast to CNNs, where convolutional layers extract features of an image in a hierarchical manner, with a limited local receptive field.
-Lightweight Transformer models have been developed, e.g., MobileViT [12], that have been applied for HW implementation [13].
 Unlike the numerical models based on neural networks mentioned above, the Tsetlin machine [10] is an ML algorithm founded on propositional logic.
-During training, parallel conjunctive clauses are defined by independent learning automata.
-Each clause represents a logical sub-pattern and will output a logical 1 if it matches an input sample, and 0 otherwise.
-The TM features a single-layer structure with highly interpretable outputs [15, 16, 17], making it well-suited for application areas such as biomedical, cybernetics, IoT, and law, where interpretability is essential.
-Furthermore, because its primary operations are Boolean, TM solutions are both low-complexity and hardware-efficient.
+Parallel conjunctive clauses are defined by independent learning automata, and each clause represents a logical sub-pattern.
+The TM features a single-layer structure with highly interpretable outputs, making it well-suited for application areas such as biomedical, cybernetics, IoT, and law.
+Furthermore, because its primary operations are of Boolean type, TM solutions are hardware-efficient.
 The generation of a TM’s output relies on summing and comparing class sums, with minimal or no need for energy-intensive multiplication operations, even during training.
 This makes TMs ideal for low-energy application-specific integrated circuit (ASIC) implementations [11].
-Additionally, due to their small resource footprint, TMs facilitate hardware solutions that support on-device training and continuous learning [19, 20].
-When compared against deep neural networks (DNNs) and other ML algorithms, the TM has demonstrated competitive performance.
-It has been tested on tabular data [14], natural language [17, 15], images [21, 22, 23] and regression [24].
-Through the works in [25, 26, 27] the convergence of the TM’s basic operators has been verified.
+Additionally, due to their small resource footprint, TMs facilitate hardware solutions that support on-device training and continuous learning [12].
 The original TM version [10] is called the vanilla TM.
 For image classification, improved test accuracy is achieved by the convolutional TM (CTM) [13].
-The CTM has obtained a peak test accuracy of 99.4% on the MNIST dataset [28].
-For the Fashion-MNIST (FMNIST) [29] and Kuzushiji-MNIST (KMNIST) [30] datasets, the peak test accuracies achieved are 91.5% and 96.31% respectively.
+The CTM achieved a peak test accuracy of 99.4% on the MNIST dataset [14].
+For the Fashion-MNIST (FMNIST) [15] and Kuzushiji-MNIST (KMNIST) [16] datasets, the peak test accuracies achieved were 91.5% and 96.31% respectively.
 Ensemble TM approaches have shown promising accuracies for more complex image datasets, such as CIFAR-10 and CIFAR-100, which are currently being investigated.
 In the case of CIFAR-10, the approach denoted TM Composites [17] has demonstrated a test accuracy of 82.8% [18].
-In this paper, we present an all-digital inference accelerator ASIC based on the coalesced TM (CoTM) [22].
-The CoTM utilizes a single clause pool, in contrast to the vanilla TM which uses one pool of clauses per class.
+In this paper, we present an inference accelerator ASIC based on the coalesced TM (CoTM) [19].
+The CoTM utilizes a single clause pool, in contrast to the vanilla TM which uses one clause pool per class.
 As we operate the CoTM with convolution, we here denote it as a convolutional CoTM (ConvCoTM).
-For configurations with few clauses, the ConvCoTM achieves better accuracy than the vanilla CTM, which makes the ConvCoTM especially attractive for HW solutions with constrained resources.
-The main motivation behind the development of this ASIC was to demonstrate the HW friendliness and power efficiency of a ConvCoTM-based accelerator. The main contributions of this paper are:
+For configurations with few clauses, the ConvCoTM achieves better accuracy than the vanilla CTM, making it attractive for resource-constrained solutions.
+Until now, only a single ASIC based on the TM has been reported [11], which operated on a simple dataset.
+We wanted to make an an accelerator that operated on a well-known dataset (MNIST) and to demonstrate that an all-digital TM-based ASIC could achieve energy-efficiency at the same level as analog in-memory-computing (IMC) solutions [20, 21, 9].
+The ASIC’s accelerator core is built on a modified version of the ConvCoTM inference core included in the FPGA solution in [12], which supported full on-device training.
+However, because of the FPGA it could not achieve ultra-low-power (ULP) operation.
+Compared to [12], several enhancements are made in the inference core of the ASIC, as described in section IV.
+In particular, there is an updated clause logic circuit with a novel feedback that reduces switching activity.
+There is also new functionality for loading a pre-trained model.
+Most important is the optimization of the inference part to achieve ULP operation.
+Although MNIST is a simple dataset, it is still widely used for benchmarking of ULP image classification ASICs [20, 21, 9].
+We therefore consider the MNIST dataset suitable for demonstrating the HW-friendliness and power-efficiency of a ConvCoTM-based accelerator, which was the main motivation for the design and implementation of this ASIC.
+The main contributions of our paper are:
   • The accelerator is the first reported manufactured ASIC based on the CoTM. It is also the first TM ASIC where convolution is utilized.
-  • The ConvCoTM accelerator core operates with an energy per classification (EPC) of 8.6 nJ.
-To the best of the authors’ knowledge, this is the lowest reported EPC for manufactured ASICs that operate on a similar test dataset and with comparable test accuracy.
+  • The accelerator is a fully digital, synchronous design, compatible with standard digital workflows.
+It operates with an energy per classification (EPC) of 8.6 nJ, which is the lowest reported for a fully digital solution running on the same dataset and with comparable test accuracy.
+Overall, it ranks as the second most energy-efficient solution.
+Other comparable solutions are of analog and mixed-signal type, utilizing IMC [20, 21, 9].
   • The accelerator has low latency, essential for real-time operation in low-power systems with strict requirements.
-  • The ConvCoTM solution is a fully digital, synchronous design compatible with standard digital workflows.
+  • For TMs operating with convolution, we introduce a novel clause logic circuit with feedback, which reduces switching of the clause’s combinational output.
+  • We describe an envisaged scaled-up ASIC that operates on the CIFAR-10 dataset, and provide estimates of its performance.
 The remainder of the paper is organized as follows: Section II summarizes related work.
@@
 It has a power consumption of 5.6 mW and obtains an accuracy of 86% when operating on the CIFAR-10 dataset.
-An architecture for a binary-weight CNN is described in [8].
-Images of size $224\times 224\times 3$ are classified at a rate of 11 frames per second (FPS).
-The technology used is 65 nm CMOS and simulations show a power consumption of 895 $\mu$W at 0.6 V.
-The accelerator ASIC in [35] is manufactured in 65 nm CMOS.
-Its classification rates for Imagenet [36] are 35/0.7 FPS for the AlexNet and VGG-16 models respectively.
-The ASIC consumes 278 mW/236 mW with corresponding EPC values of 7.9 mW/0.4 mJ.
-The successor to [35] is the work described in [37].
-Here, additional steps are taken to reduce energy consumption further.
-In particular, it reduces the movement of all data types during the various operational steps of the implemented CNN algorithm.
-A highly flexible on-chip network is utilized and sparse data are processed efficiently.
-When operating on $224\times 224\times 3$ images from Imagenet, simulations show a power consumption of 1.44 mW at 1470.6 FPS.
-The technology used is 65 nm CMOS.
-In-memory-computing (IMC) solutions can provide very high energy efficiency, as the bottleneck related to memory access is relieved.
+IMC solutions can provide very high energy-efficiency, as the bottleneck related to memory access is relieved.
 IMC has especially been utilized for analog/mixed-signal ML accelerator designs.
@@
 This accelerator achieves an EPC of 3.8 $\mu$J when operating on CIFAR-10.
-The classification rate is 237 FPS with a power consumption of 0.9 mW.
-In [9] a charge-domain IMC ternary neural network (TNN) classifier is described.
+The classification rate is 237 frames per second (FPS) with a power consumption of 0.9 mW.
+In [9] a charge-domain IMC ternary neural network (TNN) chip is described.
 The main advantage of the ternary approach is a reduction of the required operations per classification with 3.9$\times$ compared to a BNN model.
-The accelerator classifies MNIST images at a rate of 549 FPS.
-It has very low power consumption, only 96 $\mu$W, which corresponds to 0.18 $\mu$J per classification.
+The accelerator classifies MNIST images at a rate of 549 FPS, consuming 96 $\mu$W.
+This corresponds to an EPC of 0.18 $\mu$J.
 The test accuracy achieved is 97.1%.
+The accelerator chip that achieves the lowest reported EPC for MNIST classification, is reported in [20].
+It is a CNN utilizing an analog IMC architecture with time domain signal processing.
+It is implemented in 28nm CMOS and achieves an EPC of only 3.32 nJ.
 Neuromorphic approaches, utilizing spiking neural networks (SNNs) have achieved very low energy consumption.
 In [21] a mixed-signal neuromorphic SNN IC is reported, which has been manufactured in a 65 nm CMOS technology.
-It achieves a test accuracy on MNIST of 95.35%, a power consumption of 0.517 mW and an EPC of 12.92 nJ.
+It achieves a test accuracy on MNIST of 95.35%, with a power consumption of 0.517 mW and an EPC of 12.92 nJ.
 Simulation results of a 28 nm CMOS charge-domain computing SNN solution are reported in [24].
 The accelerator achieves an EPC of 15.09 nJ, a latency of 0.46 $\mu$s and a test accuracy on MNIST of 94.81%.
-In [40] simulation results of a clock-free, event-driven, mixed-signal 28 nm IMC SNN are described.
+In [25], simulation results of a clock-free, event-driven, mixed-signal 28 nm IMC SNN are described.
 On MNIST it achieves a test accuracy of 96.92% and 1.38 nJ EPC.
-For the TM-based hardware solutions, the first reported IC is described in [18].
+For TM-based HW solutions, the first reported ASIC is described in [11].
 It utilizes the vanilla TM architecture and operates on a 3-class machine learning (ML) problem.
-The IC is implemented in a 65 nm CMOS technology and supports both training and inference.
+The chip is implemented in a 65 nm CMOS technology and supports both training and inference.
 In [26], the first TM-based HW accelerator with convolution is reported.
@@
 Full on-device training is implemented.
-The FPGA solution reported in [41] is based on the vanilla TM and stores and processes only the include signals of the TA actions, utilizing the typical TM model’s high sparsity.
-It therefore represents a highly compact implementation when it comes to HW resources.
-The design operates on the MNIST dataset, with sequential processing of each clause and each literal.
-In [20] the first reported ConvCoTM-based HW solution is presented.
-It is an FPGA solution that classifies $28\times 28$ images with 10 categories, and includes complete on-device training.
+The FPGA solution in [27] is based on the vanilla TM and utilizes the sparsity of typical TM models, in combination with sequential operation.
+It represents a highly compact solution when it comes to HW resource usage.
+In [12] the first reported ConvCoTM-based HW solution is presented, of which the inference part is the basis for our ASIC as described in the Introduction section.
+It is an FPGA solution that classifies $28\times 28$ images with 10 categories, and includes full on-device training.
 For the MNIST dataset a test accuracy of 97.6% is achieved, with a classification rate of 134k samples per second, and an EPC of 13.3 $\mu$J.
-Several articles describe the design and simulation of HW solutions for TM accelerators.
-In [42, 43, 44] low-latency asynchronous TM solutions for inference and training are presented.
-A mixed-signal IMC TM solution is described in [45] with an EPC for MNIST of 13.9nJ.
+The source code for this design is available at [28].
+In [29, 30, 31] the design and simulation of low-latency asynchronous TM solutions for inference and training are presented.
+A mixed-signal IMC TM solution is described in [32] with a simulated EPC for MNIST of 13.9 nJ.
 It is an inference solution with a Boolean-to-current architecture, and resistive RAM (ReRAM) transistor cells are employed to eliminate the need for energy-hungry analog-to-digital and digital-to-analog conversions.
-Another IMC TM concept based on Y-flash cells is reported in [46].
-Based on the above literature review, developing a ConvCoTM ASIC solution is a critical step toward achieving higher energy efficiency, performance, and scalability for TM-based hardware accelerators.
-Existing FPGA implementations, while effective, are limited in energy efficiency.
-Comparatively, in-memory computing (IMC) and binary neural network (BNN)-based ASICs demonstrate significantly lower energy consumption and higher operational efficiency, underscoring the potential benefits of an ASIC approach.
-By leveraging a standard digital design flow and CMOS technology, an ASIC solution for ConvCoTM could deliver real-time, power-efficient classification.
-This would position ConvCoTM as a competitive alternative to CNNs and BNNs in resource-constrained environments.
+Another IMC TM design concept based on Y-flash cells, is reported in [33].
 III TM and CoTM Background
@@
 $L=[l_{0},l_{1},\dots,l_{2o-1}]\\ =[x_{0},\dots,x_{o-1},\neg x_{0},\dots,\neg x_{o-1}]$ (1)
-The Tsetlin Automaton (TA) of two-action type [47] constitutes the basic learning element in a TM.
-The TA will either include or exclude a literal in a conjunctive clause, depending on its state.
-One team of TAs is applied per clause, and the different TA actions are obtained during the training process which involves feedback mechanisms to each TA [14].
-For an inference-only solution, as reported in this paper, we need only the TA action signals from the trained model, not the complete TAs.
-In a TM the number of clauses, $n$, is a user specified integer. The output of a single clause, $c_{j}$, where $j\in\{0,\ldots,n-1\}$, is given by
+A TA of two-action type [34] will either include or exclude a literal in a conjunctive clause, depending on its state.
+One team of TAs is applied per clause, and the different TA actions are obtained during the training process, which involves feedback mechanisms to each TA [10].
+For an inference-only solution, as reported in this paper, only the TA action signals from the trained model are needed, not the complete TAs.
+In a TM, the number of clauses, $n$, is a user specified integer. The output of a single clause, $c_{j}$, where $j\in\{0,\ldots,n-1\}$, is given by
 $c_{j}=\bigwedge_{k\in I_{j}}l_{k},$ (2)
-where $l_{k}$ is the literal with index k. k belongs to $I_{j}\subseteq\{0,....,2o-1\}$, where $I_{j}$ is the set of literal indexes for clause $j$ where the TA action is include [14].
+where $l_{k}$ is the literal with index k.
+$k$ belongs to $I_{j}\subseteq\{0,\dots,2o-1\}$, where $I_{j}$ is the set of indices of the literals in clause $j$ for which the corresponding TAs take the action include.
 III-B CoTM
@@
 The clause weights in Eq.
-(3) obtain their values during training and can have both positive and negative polarity.
-It should be noted that no multiplications are needed to implement Eq.
-(3), only additions, as a clause can output only 0 or 1.
+(3) can have both positive and negative polarity, and obtain their values during training.
+No multiplications are needed to implement Eq.
+(3), only additions, as a clause can output either 0 or 1.
 The predicted class, $\hat{y}$, is determined by the argmax operator, as given by Eq.
@@
 III-C CoTM and Convolution
-One can improve the image classification test accuracy by applying convolution [21] to the CoTM.
-In addition, the number of features that need to be processed simultaneously, is lowered with convolution, which again reduces the model size and the complexity of a design.
-However, the benefits of convolution come at the expense of increased processing time, which is due to the need to sequentially process the different patches of the input sample.
+It has been shown that applying convolution to the TM can improve the image classification test accuracy [13].
+Convolution also reduces the number of features that need to be processed simultaneously, which in turn reduces model size and design complexity.
+However, the benefits of convolution come at the cost of increased processing time, due to the need to sequentially process the various patches of the input sample.
 For image classification, the input data samples to a ConvCoTM, are a set of images, each of dimensions $X\times Y$, with $Z$ channels.
@@
 The convolution window will be evaluated $B=B_{X}\times B_{Y}$ times across the image, where $B_{X}=1+(X-W_{X})/d_{x}$, $B_{Y}=1+(Y-W_{Y})/d_{y}$, and $d_{x}$ and $d_{y}$ are the stride values of the convolution window in the $X$ and $Y$ directions respectively [13].
-For each window position, the number of Boolean features is given by
+For each window position, the number of Boolean features generated per patch is given by
 $N_{F}=W_{X}\times W_{Y}\times Z\times U+(Y-W_{Y})+(X-W_{X}).$ (5)
@@
 (1), the ConvCoTM also takes as input the negated versions of the features to form the literals per patch.
+Figure 1: ConvCoTM accelerator ASIC block diagram.
 III-D Datasets and Booleanization
@@
 For MNIST, they are converted into Boolean variables (i.e., booleanized) with simple thresholding, where pixel values larger than 75 are replaced with 1, and 0 otherwise [13].
-For FMNIST and KMNIST, the booleanization is performed with an adaptive Gaussian thresholding procedure [21].
-Fig.
-1 shows one example of an original image and the corresponding booleanized image from each of the three datasets.
-For the booleanized images the color white represents a logical 1.
-Figure 1: Examples of MNIST, FMNIST and KMNIST image booleanization.
-For the booleanization, we use one bit per pixel, i.e., $U=1$ for all three datasets.
-The convolution window applied has dimensions $W_{X}=W_{Y}=10$, and a step size of 1 is used in both directions.
-The number of feature bits per patch is therefore 100 bits from the convolution window plus 36 bits for the appended patch position information, ref.
+For FMNIST and KMNIST, an adaptive Gaussian thresholding procedure [13] is applied.
+We use one bit per pixel, i.e., $U=1$ for all three datasets.
+The convolution window has dimensions $W_{X}=W_{Y}=10$, and a step size of one is used in both directions.
+For each patch, we have 100 feature bits from the convolution window plus 36 bits for the appended patch position information, ref.
 Eq.
 (5).
-The number of literals for each patch is 272, i.e., two times the number of features.
+The number of literals for each patch is therefore 272.
 III-E TM and Convolution - Inference
 During inference, each clause in a ConvCoTM is evaluated $B$ times per image.
-After finishing the convolution, a clause will output 1 if it has recognized a pattern at least once in any of the $B$ patches for a given image [21].
-The clause output is given by the sequential OR-function in Eq.
-(6), where $c_{j}$ is the $j$’th clause of the ConvCoTM, and $c_{j}^{b}$ is the $b$’th output of this clause obtained during the window sliding (patch generation).
-$c_{j}=\bigvee_{b=0}^{B-1}c_{j}^{b},$ (6)
+After finishing the patch generation, a clause will output 1 if it has recognized a pattern at least once in any of the $B$ patches for a given image [13].
+The clause output is given by Eq.
+(6) which we call the sequential OR-function.
+Here $c_{j}$ is the $j$’th clause of the ConvCoTM, and $c_{j}^{b}$ is the $b$’th output of this clause obtained during the patch generation:
+$c_{j}=\bigvee_{b=0}^{B-1}c_{j}^{b}.$ (6)
 Class sum generation and class prediction are similar to the case without convolution, and follow Eq. (3) and Eq. (4).
@@
 The block diagram of the ConvCoTM accelerator is shown in Fig.
-2.
+1.
 The accelerator includes all circuitry required for performing inference, and only minimal interaction with the system processor is required.
 The three main modules are i) a data interface for transfer of model parameters and image samples, ii) a module for model storage and iii) an inference core that performs classification of an image.
-Figure 2: ConvCoTM accelerator ASIC block diagram.
-The main configuration parameter of a TM solution is the number of clauses.
-Due to chip area restrictions, we decided to implement a configuration with 128 clauses.
 The following subsections describe the different sub-modules and the accelerator’s operation in detail.
@@
 All other operations are performed by the accelerator.
-The data interface towards the external system processor, is an 8-bit parallel type inspired by the ARM AXI Stream interface [50].
+The 8-bit parallel data interface towards the external system processor, is inspired by the ARM AXI Stream interface [36].
 It is used for transferring model data to the accelerator during load model mode, and for transferring images in inference mode.
 When a classification operation is finished, the accelerator outputs an interrupt.
-In addition, it outputs a separate 8-bit signal consisting of the predicted class (4 bits) together with the true label of the sample (4 bits).
+In addition, it outputs a separate 8-bit signal consisting of the predicted class of the sample together with the true label (if provided).
 IV-B ConvCoTM Model Registers
@@
 IV-C Image Data Buffer and Patch Generation
-A data buffer is included with room for two complete $28\times 28$ booleanized images.
-While one image is being processed, another image can be transferred from the external system processor to the other image buffer, thereby improving throughput.
-We denote this as the continuous classification mode.
-A complete booleanized image requires $28\times 28/8=98$ bytes.
-In addition, one byte is allocated for the image label (0 to 9).
+The accelerator includes a data buffer with room for two complete $28\times 28$ booleanized images.
+While one image is being processed, another image can be transferred from the external system processor to the other image register, thereby improving throughput.
+We denote this as continuous classification mode.
+The size of a complete booleanized image is $28\times 28/8$ bytes $=98$ bytes.
+One additional byte is allocated for the image label (0 to 9).
 The patch generation module consists of a register with 10 rows, each with 28 DFFs.
 At the start of processing a new sample, the first 10 rows of an image are loaded into the register, and the convolution window starts at the leftmost position, which has coordinates ($0,0$).
-During the convolution, the window is moved to the right with increments of one.
+During the patch generation, the window is moved to the right with increments of one.
 As described in Subsection III-C, the position of the window is thermometer encoded.
 There are 19 different $x$-coordinates for the window, requiring 18 bits.
-When the window has been evaluated at the rightmost $x$-position (0,18), the contents of all 10 rows are shifted upwards with one step, and the next image datarow is loaded into the lowermost row.
-Then the window starts again at the leftmost position, now with patch coordinates (1,0), and moves to the right in steps of one.
+When the window has been evaluated at the rightmost $x$-position ($0,18$), the contents of all 10 rows are shifted upwards with one step, and the next image datarow is loaded into the lowermost register row.
+The convolution window then starts again at the leftmost position, now with patch coordinates ($1,0$).
 This procedure is repeated until all 28 datarows of the image have been processed.
-Also for the $x$-coordinate, there are 19 different positions.
-During a complete convolution, $19\times 19=361$ patches are generated.
+There are 19 different positions also for the $y$-coordinate, and during a complete patch generation phase, $19\times 19=361$ patches are generated.
 IV-D Clause Pool
 The ConvCoTM is configured with 128 clauses that operate in parallel.
-Each of the 128 clauses takes the literals, $[l_{0},l_{1},\dots,l_{271}]$, as input, as well as their corresponding TA action signals, $[i_{0},i_{1},\dots,i_{271}]$, from the model register.
+Each clause takes the literals, $[l_{0},l_{1},\dots,l_{271}]$, as input, as well as their corresponding TA action signals, $[i_{j,0},i_{j,1},\dots,i_{j,271}]$, $j\in\{0,\ldots,127\}$, from the model register.
 Fig.
@@
 2 constitutes a single-bit register for clause $j$.
-During the convolution, its content, $c_{j}$, is ORed with the combinational clause value, $c_{j}^{b}$, of the current patch, thus implementing the sequential OR function in Eq.
+Its content, $c_{j}$, is ORed with the combinational clause value, $c_{j}^{b}$, of the current patch, during the patch generation.
+This implements the sequential OR function in Eq.
 (6).
-Before starting a new convolution (patch generation), the DFF is reset.
+The DFF is reset before starting a new convolution process.
 The signal $c_{j}$ is also fed back to the input of the OR-gates, as there is no need for the combinational logic to evaluate a clause for more patches if $c_{j}$ already is high.
-This feedback mechanism, hereinafter denoted the clause switching reduction feedback (CSRF), reduces the digital switching in the combinational part of the clause logic.
-Simulations on MNIST image classification with CSRF enabled, showed an average of 50% reduction in the toggling rate of $c_{j}^{b}$ for each clause.
+This feedback, hereinafter denoted the clause switching reduction feedback (CSRF), reduces the switching in the combinational part of the clause logic.
+Simulations of MNIST classification utilizing CSRF, showed an average of 50% reduction in the toggling rate of $c_{j}^{b}$ for each clause.
 The accelerator ASIC has a dedicated pin that can enable or disable the CSRF.
@@
 In this case, the combinational clause logic that generates the signal $c_{j}^{b}$ in  Fig.
-3, would have to be replicated for each convolution window.
+2, would have to be replicated for each convolution window.
 IV-E Inference
-The inference operation starts with loading an image into the data buffer, followed by a convolution process, as described in Subsection IV-C.
+The inference operation starts with loading an image into the data buffer, followed by sequential generation of the patches, as described in Subsection IV-C.
 When finished, each clause value, $c_{j}$, is weighted, for each class, according to Eq.
 (3).
-All ten class sums are generated in parallel, by adders configured in a tree structure and utilizing a three-stage pipeline.
+All ten class sums are generated in parallel, by adders configured in a tree-structure and utilizing a three-stage pipeline.
 Finally, an argmax module [11] selects the class with the highest class sum, as the predicted class, $\hat{y}$.
@@
 Figure 3: Simplified accelerator state diagram.
-The accelerator’s latency, measured from when the system processor initiates the first byte transfer of the image, until a prediction result is available, is 471 clock cycles.
-This includes 99 clock cycles for transferring the 98 image bytes and the label byte, and 372 clock cycles for the patch generation and class prediction.
-The accelerator can also operate in continuous mode, where a new sample is loaded into the image buffer while the current sample is being processed, see Subsection IV-C.
+The accelerator’s latency is 471 clock cycles.
+This is measured from when the system processor initiates the first byte transfer of the image, until a prediction is available.
+It includes 99 clock cycles for transferring the 98 image bytes and the label byte, and 372 clock cycles for the patch generation, class summation and class prediction.
+When the accelerator operates in continuous mode, a new sample is loaded into the image buffer while the current sample is being processed, see Subsection IV-C.
 This increases the throughput, as image samples can be processed every 372’th clock cycle.
@@
 IV-F Clock Domains
-The key to reducing the digital switching power of the accelerator ASIC lies at the top level of the ASIC architecture.
-Two separate clock domains are utilized, each with its own dedicated clock pin, assigned to i) the model part and ii) the inference core.
-When a model has been transferred from the system processor to the ASIC, the TA action signals and the weights are available in registers which do not need further clocking.
-By stopping the clock to the model part when operating in inference mode, the digital switching power is reduced significantly, as the model part constitutes approximately 90% of the accelerator’s DFFs.
+The key to reducing the digital switching power of the accelerator, lies at the top level of the ASIC architecture.
+Two separate clock domains are utilized, each with its own dedicated clock pin, assigned to the model part and the inference core, respectively.
+When a model has been transferred from the system processor to the ASIC, the TA action signals and the weights are available in registers that do not need further clocking.
+By stopping the clock to the model part, when operating in inference mode, the digital switching power is reduced significantly, as the model part constitutes approximately 90% of the accelerator’s DFFs.
 To further reduce power consumption, standard clock gating is applied for the inference module.
-This ensures, for instance, that the pipelined registers associated with the class sum generation are enabled and clocked only for 4 clock cycles per classification cycle.
+This ensures, for instance, that the pipelined registers associated with the class sum generation are enabled and clocked only for four clock cycles per classification phase.
 The clock-gating can be enabled/disabled by an external pin.
+Figure 4: Photo and layout plot of the ConvCoTM accelerator chip.
 V Implementation and Measurement Results
-The ConvCoTM accelerator was designed in VHDL and first implemented and verified on a Xilinx ZCU104 FPGA development board.
-The ASIC was then implemented in a 65 nm low-leakage CMOS technology from UMC, and packaged in a JLCC44 ceramic package.
+The ConvCoTM accelerator was designed in VHDL and implemented in a 65 nm low-leakage CMOS technology from UMC.
+A JLCC44 ceramic package housed the chip.
 Fig.
 4 shows a photograph and a layout plot of the ASIC.
-The accelerator’s core area is 2.7 mm2.
-Figure 5: Photo and layout plot of the ConvCoTM accelerator chip.
-A test PCB was developed for the ASIC.
-This board was connected to a Digilent Zybo Z7-20 development board equipped with a Xilinx Zynq XC7Z020 FPGA, as shown in Fig.
-6.
-One of the FPGA’s ARM9 cores was configured as system processor, and the FPGA board provided clocks, model data, image data samples and control signals to the ASIC accelerator.
-The processor also read back accelerator status/interrupt signals and the predicted/true class per sample.
+The accelerator’s core area is approximately 2.7 mm2.
+The VHDL source code for the ASIC implementation (used for initial FPGA verification) is available at a GitHub repository [37].
+A printed circuit board (PCB) was developed for testing and characterization of the ASIC.
+The ASIC PCB was connected to a Digilent Zybo Z7-20 development board, equipped with a Xilinx Zynq XC7Z020 FPGA, as shown in Fig.
+5.
+One of the FPGA’s ARM9 cores was configured as system processor, and the FPGA board provided clocks, model data, image data samples and control signals to the ASIC PCB.
+The processor also read back accelerator status signals, and the predicted class per sample.
 To interface with the FPGA board, the ASIC’s input/output (IO) ports operated at 3.3 V supply voltage.
@@
 Maximum/minimum limits were set on the clause weights to fit with the allocated 8-bits per weight in the accelerator.
-For simplicity, the images from the test dataset were preprocessed (booleanized), as described in subsection III-D, and included in the FPGA processor’s test program.
+For simplicity, the images from the test datasets were preprocessed (booleanized), as described in Subsection III-D, and included in the FPGA processor’s test program.
 Model data and image samples were applied from the FPGA to the accelerator via the AXI Stream interface.
-We utilized DMA functionality on the FPGA to enable fast data transfer.
+We utilized direct memory access (DMA) functionality on the FPGA to enable fast data transfer.
 The maximum classification rate achieved, including system processor timing overhead, was 60.3k images per second in continuous mode, with an accelerator clock frequency of 27.8MHz.
-The speed limiting factor was the electrical interface between the FPGA and the ASIC.
-The latency when classifying a single image was 25.4 $\mu$s, including the transfer time of the image to the accelerator as well as timing overhead in the system processor.
+The maximum operating frequency was limited by the general purpose IOs (GPIOs) on the FPGA board.
+When classifying a single image, the latency was 25.4 $\mu$s, including the transfer time of the image to the accelerator as well as timing overhead in the system processor.
 The achieved test accuracies of the datasets MNIST, FMNIST and KMNIST were 97.42%, 84.54% and 82.55% respectively.
 These results are independent of the accelerator clock frequency and are exactly in accordance with the performance of the models obtained from the SW simulations.
-In a real application, the ConvCoTM design would typically be included as a processor peripheral module in an SoC.
-Therefore, we considered the power consumption of the accelerator core as the most interesting performance power parameter, and our measurements excluded energy consumed by the digital 3.3 V IOs of the ASIC.
-Table I summarizes the characteristics and performance of the accelerator.
 TABLE I: The ConvCoTM accelerator ASIC’s main characteristics and performance
@@
 9.6 nJ b,d
-Latency (single image classi-, 25.4 $\mu$s a
+Latency (single image classi- 25.4 $\mu$s a
 fication including data transfer) 0.66 ms b
@@
 This corresponds to an EPC of 19.1 nJ.
-With a supply voltage of 0.82 V, the power consumption was 0.52 mW and the corresponding EPC was 8.6 nJ.
-With 1.2 V core supply voltage and a clock frequency of 1 MHz, the achieved classification rate was 2.27k images/second at a power consumption of 81 $\mu$W.
+With a supply voltage of 0.82 V, the accelerator core’s power consumption was 0.52 mW and the corresponding EPC was 8.6 nJ.
+With 1.2 V core supply voltage and a clock frequency of 1 MHz, the achieved classification rate was 2.27k FPS at a power consumption of 81 $\mu$W.
+Table I summarizes the characteristics and performance of the accelerator.
+In a real application, the ConvCoTM design would typically be included as a peripheral module to a microcontroller in an SoC.
+Therefore, we considered the power consumption of the accelerator core as the most interesting power parameter.
+In inference mode, the digital 3.3 V IOs, which were required for interfacing to the FPGA board, had a power consumption of 0.76 mW, at a clock frequency of 27.8 MHz.
 All power and energy measurement results of the ASIC are average values, and were performed with clock-gating and CSRF enabled (as shown in Fig.
-3).
+2).
 When operating at 27.8 MHz, clock-gating reduced the power consumption by approximately 60%, while the CSRF alone provided less than 1% power reduction.
 VI Discussions
-We consider the ConvCoTM solution with 128 clauses as a small configuration.
-In light of this, we are satisfied with the obtained test accuracy which is comparable with other low-complexity and low-power solutions.
-Table II shows the performance and properties of state-of-the-art low-power IC implementations for image classification.
-Inspired by the TinyML [52] benchmarks for embedded systems we specify the classification rate, the peak power consumption and the EPC.
+We consider the ConvCoTM accelerator with 128 clauses as a small configuration.
+In light of this, we are satisfied with the obtained test accuracy which is comparable with other low-complexity and ULP solutions.
+Table II shows the performance and properties of state-of-the-art ULP IC implementations for classification of MNIST and CIFAR-10 datasets.
+Inspired by the TinyML [39] benchmarks for embedded systems, we specify the classification rate, the peak power consumption and the EPC.
 These are system parameters which are independent of the accelerator model and architecture [40].
 However, when used for comparing solutions, one must be careful to specify the ML task performed.
-The ICs shown in Table II operate on images and datasets that range from low to high complexity.
-Accelerators that operate on CIFAR-10 and other more complex images, typically have lower classification rates and higher EPC compared to accelerators that operate on $28\times 28$ images.
+The CIFAR-10 dataset is significantly more challenging than MNIST.
+Therefore, accelerators that operate on CIFAR-10, have lower classification rates and higher EPC compared to accelerators that operate on MNIST.
 TABLE II: Comparison of the ConvCoTM IC accelerator with previous works.
-This work TCAS-II’23 Yejun [38] JSSC’23       Yang [9] TCAS-I’20 Mauro [6] JSSC’21       Knag [7] TCAS-I’20 Bankman [5] JSSC’17 Chen [35] JETCAS’19 Chen [37]
-Technology 65 nm CMOS 65 nm CMOS 40 nm CMOS 22 nm FD-SOI 10 nm FinFET 28 nm CMOS 65 nm CMOS 65 nm CMOS
-Active area 2.7 mm2 0.57 mm2 0.98 mm2 2.3 mm2 0.39 mm2 4.6 mm2 12.3 mm2 Not stated
-Algorithm ConvCoTM SNN Ternary CNN BNN BNN BNN CNN Sparse DNN
-Design type Digital Neuromorphic mixed-signal IMC, mixed-signal Digital (SoC) Digital IMC, mixed-signal Digital Digital
-Image dataset MNIST, FMNIST, KMNIST MNIST MNIST CIFAR-10 CIFAR-10 CIFAR-10 Imagenet Imagenet
-Measured test accuracy 97.42%, 84.54%, 82.55% 95.35% 97.1% 99% of nominal accuracy 86% 86% Not stated 79.6% i           79.7% j
-Classifications per second 60.3k a           2.27k b 233k (1.2 V)           40k (0.7 V) 549 15.4 Not stated 237 35 g           0.7 h 278 f,i           1470 f,j
-Latency e 25.4 $\mu$s a           0.66 ms b Not stated Not stated Not stated Not stated Not stated 115.3 ms Not stated
-Inference power 1.15 mW a,c 0.52 mW a,d 81 $\mu$b,c 21 $\mu$b,d 9.37 mW (1.2 V) 0.517 mW (0.7 V) 96 $\mu$W 674 $\mu$W 5.6 mW (13 MHz) 607 mW (622 MHz) 0.9 mW 278 mW g 236 mW h Not stated
-Energy per classification (EPC) 19.1 nJ a,c           8.6 nJ a,d           35.3 nJ b,c           9.6 nJ b,d 40.17 nJ (1.2 V) 12.92 nJ (0.7 V) 0.18 $\mu$J 43.8 $\mu$J Not stated 3.8 $\mu$J 7.9 mJ g           337 mJ h 1.5 mJ f,i 0.4 mJ f,j
+This work TCAS-I´25 Zhao [20] TCAS-II’23 Yejun [21] JSSC’23       Yang [9] TCAS-I’20 Mauro [6] JSSC’21       Knag [7] TCAS-I’20 Bankman [5]
+Technology 65 nm CMOS 28 nm CMOS 65 nm CMOS 40 nm CMOS 22 nm FD-SOI 10 nm FinFET 28 nm CMOS
+Active area 2.7 mm2 0.261 mm2 0.57 mm2 0.98 mm2 2.3 mm2 0.39 mm2 4.6 mm2
+Algorithm ConvCoTM CNN SNN Ternary CNN BNN BNN BNN
+Design type Digital Analog, time domain Neuromorphic mixed-signal IMC, mixed-signal Digital (SoC) Digital IMC, mixed-signal
+Image dataset MNIST, FMNIST, KMNIST MNIST MNIST MNIST CIFAR-10 CIFAR-10 CIFAR-10
+Measured test accuracy 97.42%, 84.54%, 82.55% 97.9% 95.35% 97.1% 99% of nominal accuracy 86% 86%
+Classifications per second 60.3k a           2.27k b 3508 233k (1.2 V)           40k (0.7 V) 549 15.4 Not stated 237
+Latency e 25.4 $\mu$s a           0.66 ms b Not stated Not stated Not stated Not stated Not stated Not stated
+Inference power 1.15 mW a,c 0.52 mW a,d 81 $\mu$b,c 21 $\mu$b,d 11.6 $\mu$W 9.37 mW (1.2 V) 0.517 mW (0.7 V) 96 $\mu$W 674 $\mu$f 5.6 mW (13 MHz) 607 mW (622 MHz) 0.9 mW
+Energy per classification (EPC) 19.1 nJ a,c           8.6 nJ a,d           35.3 nJ b,c           9.6 nJ b,d 3.32 nJ 40.17 nJ (1.2 V) 12.92 nJ (0.7 V) 0.18 $\mu$J 43.8 $\mu$J Not stated 3.8 $\mu$J
 a 27.8 MHz clock frequency.
@@
 d vdd=0.82 V.e For processing of a single image sample, including data transfer.
-f Simulated results.
-g AlexNet.
-h VGG-16.
-i Sparse AlexNet.
-j Mobilenet.
-The mixed-signal IMC-based solution in [9] operates with a very low peak power consumption of 96 $\mu$W and consumes 0.18 $\mu$J per classification on MNIST with a test accuracy of  97.1%.
-Its classification rate is 549.
-In [38] a mixed-signal SNN IC accelerator achieves a very low EPC of 12.92 nJ with 0.7 V supply voltage.
-Operating at 1.2 V the EPC is 40.17 nJ.
+f Full SoC.
+The mixed-signal IMC-based solution in [9] operates with a peak power consumption of 96 $\mu$W and consumes 0.18 $\mu$J per classification on MNIST, with a test accuracy of  97.1%.
+In [21] a mixed-signal SNN IC accelerator achieves an EPC of 12.92 nJ with 0.7 V supply voltage.
 It utilizes a time-domain echo state network (ESN) [41] approach and thus requires non-standard interfacing with the surrounding system’s resources.
-It obtains a test accuracy on MNIST of 95.35%.
-To the authors’ knowledge, the ConvCoTM accelerator core’s EPC of 8.6 nJ at 0.82 V operating voltage is the lowest reported for a manufactured ASIC that classifies $28\times 28$ greyscale images and which has comparable test accuracy.
-At 1.2 V operating voltage it obtains an EPC of 19.1 nJ, which is also very low.
-As the ConvCoTM ASIC is fully digital, the classification rate and thereby also the power consumption, can easily be scaled by adjusting the clock frequency.
-E.g, with a clock frequency of 1 MHz, the power consumption achieved is only 81 $\mu$W.
-As the accelerator is all-digital, the test accuracy is independent of the clock frequency.
-Several analog/mixed-signal IC accelerators obtain very low power consumption [38, 9, 5].
-Such solutions are susceptible to variations in process parameters, supply voltage and operating temperature (PVT), which can affect the performance.
-All-digital solutions avoid such non-ideal effects [6, 7, 8].
-In addition, they are easier to integrate in larger SoCs, and can be scaled and re-targeted to other technologies.
-This applies for CNNs as well as for TMs.
+The test accuracy achieved on MNIST is 95.35%.
+The 28 nm CMOS CNN-based accelerator in [20] utilizes an IMC architecture with analog and time-domain signal processing.
+It achieves an EPC of 3.32 nJ, which is the lowest reported for MNIST.
+Our ConvCoTM accelerator ASIC achieves an EPC of 8.6 nJ at 0.82 V operating voltage.
+This is the second lowest EPC reported, for a manufactured ASIC that operates on MNIST.
+As the ConvCoTM ASIC is fully digital, the classification rate and thereby also the peak power consumption, can be scaled by adjusting the clock frequency.
+At a clock frequency of 1 MHz, the power consumption is 81 $\mu$W.
+Most ULP accelerators are based on analog techniques [20, 21, 9, 5].
+The test accuracy of such solutions is susceptible to variations in process parameters, supply voltage and operating temperature (PVT).
+All-digital solutions, like our ConvCoTM ASIC and [6, 7, 8], avoid such non-ideal effects.
+In addition, these accelerators are easier to integrate in larger SoCs, and can more easily be scaled and re-targeted to other technologies.
+This applies for CNNs as well as TMs.
 In the ConvCoTM accelerator, the combinational clause logic draws only a small amount of energy compared to the clock tree of the inference-core DFFs.
-This is the main reason why the CSRF technique, see Subsection IV-D, did not provide a greater reduction in EPC.
+This is the main reason why the CSRF technique (see Subsection IV-D) did not provide a greater reduction in EPC.
 Other convolutional TM architectures, that operate at higher clock frequencies, process more literals in parallel, and utilize sparsity in TM models, can potentially benefit more from the CSRF technique.
-The ConvCoTM accelerator’s architecture described in this paper is highly suited for scaling.
-It can be modified to operate on larger images including 3-channel (RGB) images, e.g., based on the TM Composites architecture [31, 32].
-From the very low energy consumption obtained for our ConvCoTM accelerator, we envisage that a TM Composites-based solution will perform well compared to previous works.
-VII Conclusion
+VII Scaling
+The ConvCoTM accelerator’s architecture described in this paper, is highly suited for scaling, and can be modified to operate on larger and more complex images.
+There is active research on how to apply the TM algorithm for larger and more complex images.
+The TM version that currently achieves the best test accuracy on CIFAR-10 is TM Composites [17, 18].
+Here, several TM Specialists are applied to an image sample.
+Examples of specializations are different booleanization techniques and convolution window sizes.
+After an image has been processed, the class sums from each TM Specialist are normalized, followed by summation per class.
+Finally, an argmax operator is applied on the composite class sums for output prediction.
+Table III shows characteristics and performance estimates for an envisaged TM Composites ASIC design, intended for inference operation on CIFAR-10 images.
+It is assumed that one configurable TM module is implemented in the ASIC, and models for four different TM Specialist configurations are applied in sequence, through transferring a new model from on-chip ultra-low-power RAM per configuration.
+Booleanization of the images, for the different TM Specialists, is assumed to be performed by dedicated logic in the HW accelerator.
+We assume a timing-optimized data interface, with line buffers [42], allowing the patch generation to start as soon as a sufficient number of image lines have been transferred to the accelerator.
+For the TM Specialists, we assume an average number of literals per patch of 1000, and by utilizing the technique described in [43], the number of literals is limited to 16 per clause.
+For the TA actions, the model for each TM Specialist therefore needs to include 16 ten-bit literal addresses per clause.
+We assume the same 65 nm CMOS technology is used as for the current ConvCoTM ASIC.
+The chip area for the CIFAR-10 accelerator is estimated by multiplying the chip area of the current ASIC (2.7 mm2) with the ratio, $R$, given by the model size of one TM Specialist (32.5 kilobyte) divided by the model size of the ASIC in this paper (5.6 kilobyte).
+We therefore get $R=32.5/5.6\approx 5.8$.
+This is a reasonable assumption because the model storage in registers (for the active TM Specialist) and the clause logic, dominate the chip area.
+We further assume 2 mm2 in additional chip area for booleanization logic, additional adders for class sum generation, and RAM modules for model storage.
+We estimate that each TM Specialist will need approximately 1000 clock cycles for the processing per sample, including booleanization.
+In addition, before a TM Specialist can start operating, its model has to be loaded from on-chip RAM.
+We assume the model is organized in memory such that a transfer rate of 32 bytes per clock cycle is enabled.
+Thus, we will need approximately 1020 clock cycles for a complete model transfer.
+A single TM Specialist will therefore require in total about 2020 clock cycles to process a sample, and with four TM Specialists we will need approximately 8080 clock cycles for a complete classification.
+With a system clock of 27.8 MHz (as for the current ConvCoTM ASIC), the classification rate would be $808027.8\times 10^{6}/$ 8080 FPS $\approx 3440$ FPS.
+An estimate of the power consumption is found by multiplying the power consumption of the current ASIC’s accelerator core, operating at 27.8 MHz (1.15 mW), with the factor $R$.
+During model loading and booleanization, we assume power consumption similar to that during inference.
+An estimate of the energy per classification is then $R\times 1.15$ mW/(3440 FPS) $\approx 1.9~{}\mu$J/frame.
+TABLE III: Envisaged ConvCoTM inference ASIC Configuration and Performance for operation on the CIFAR-10 Dataset, based on the TM Composites architecture.
+Parameter Value Comment
+Number of TM 4 $4\times 4$ color thermometer,
+specialists $3\times 3$ color thermometer,
+$32\times 32$ histogram of gradients,
+$10\times 10$ adaptive thresholding.
+Number of clauses 1000
+Number of included 16 We assume 1000 literals
+literals per clause per patch, and literal limitation
+based on [43].
+Model size: TA actions 20 1000 clauses, each with 16
+per TM Specialist kilobyte literals. The address
+length per literal is 10 bits.
+Model size: Weights 12.5k 10 classes, 1000 clauses
+per TM Specialist kilobyte and 10-bit weights.
+Complete model size 130 For four TM Specialists.
+kilobyte
+Classification rate 3440 FPS Sequential operation of the
+TM Specialists, 27.8 MHz
+system clock, complete model
+stored on-chip, without
+system processor overhead.
+Test accuracy 79% Estimate for CIFAR-10.
+Accelerator core 17.7 mm2 65 nm CMOS.
+Accelerator core 6.7 mW Inference mode,
+power consumption 27.8 MHz system clock.
+Accelerator core 1.9 $\mu$J Inference mode,
+energy per 27.8 MHz system clock.
+classification.
+VIII Conclusion
 In this work, an ASIC implementation of a ConvCoTM accelerator for inference is presented.
@@
 With a clock frequency of 27.8 MHz, the accelerator’s throughput is 60.3k images per second, and with an operating voltage of 0.82 V, the energy consumed by the accelerator core per classification is 8.6 nJ.
-To the authors’ best knowledge this is lower than for previously reported fabricated ICs for the same classification task and for similar test accuracy.
-The ASIC demonstrates the simplicity, hardware-friendliness and power-efficiency of TM solutions.
-We believe the TM represents an attractive ML solution for low-power edge nodes in IoT systems.
-Furthermore, in a TM ASIC, full on-device training capabilities can be included, enabling adaptive edge solutions.
+We believe the TM represents an attractive ML solution for low-power edge nodes in IoT systems for various workloads.
+The ASIC demonstrates the simplicity, hardware-friendliness and power-efficiency of all-digital TM solutions.
 Acknowledgments

### v3

@@
 Tsetlin machine, Accelerator, Machine learning, Image classification.
-publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
+publicationid: pubid: Copyright 2025 IEEE. Personal use is permitted, but republication/redistribution requires IEEE permission.
 I Introduction
@@
 CNNs comprise many layers of multiply-and-accumulate (MAC) operations, each with their associated memory resources.
-As such, they require large number of micro-architectural operations for each classification and have significant energy footprints [3].
+As such, they require a large number of micro-architectural operations for each classification and have significant energy footprints [3].
 The complexity and energy consumption can be reduced by compromising the data precision [4].
@@
 The TM features a single-layer structure with highly interpretable outputs, making it well-suited for application areas such as biomedical, cybernetics, IoT, and law.
-Furthermore, because its primary operations are of Boolean type, TM solutions are hardware-efficient.
+Furthermore, because its primary operations are of Boolean type, TM solutions are HW-friendly.
 The generation of a TM’s output relies on summing and comparing class sums, with minimal or no need for energy-intensive multiplication operations, even during training.
@@
 For configurations with few clauses, the ConvCoTM achieves better accuracy than the vanilla CTM, making it attractive for resource-constrained solutions.
-Until now, only a single ASIC based on the TM has been reported [11], which operated on a simple dataset.
+Until now, only a single ASIC based on the TM has been reported [11], and it operated on a simple dataset.
 We wanted to make an an accelerator that operated on a well-known dataset (MNIST) and to demonstrate that an all-digital TM-based ASIC could achieve energy-efficiency at the same level as analog in-memory-computing (IMC) solutions [20, 21, 9].
-The ASIC’s accelerator core is built on a modified version of the ConvCoTM inference core included in the FPGA solution in [12], which supported full on-device training.
-However, because of the FPGA it could not achieve ultra-low-power (ULP) operation.
-Compared to [12], several enhancements are made in the inference core of the ASIC, as described in section IV.
+The ConvCoTM FPGA solution in [12] demonstrated inference and full-on device training.
+However, ultra-low-power (ULP) operation could not be achieved with the FPGA.
+As full on-device TM-based training had been proved feasible in [12], we wanted to design an inference-only ASIC solution, optimized for low-power operation.
+The ASIC accelerator’s inference core is built on modules included in [12], with several enhancements as described in section IV.
 In particular, there is an updated clause logic circuit with a novel feedback that reduces switching activity.
@@
   • The accelerator has low latency, essential for real-time operation in low-power systems with strict requirements.
-  • For TMs operating with convolution, we introduce a novel clause logic circuit with feedback, which reduces switching of the clause’s combinational output.
+  • For TMs operating with convolution, we introduce a novel clause logic circuit with feedback, which reduces switching of a clause’s combinational output.
   • We describe an envisaged scaled-up ASIC that operates on the CIFAR-10 dataset, and provide estimates of its performance.
@@
 Section III covers TM and CoTM background, including convolution operation.
-Section IV details the ASIC’s architecture and building blocks.
-Measurement results are presented in Section V, followed by a discussion in Section VI before we conclude the paper.
+Section IV details the ASIC’s architecture and building blocks, and measurement results are presented in Section V.
+In Section VI we discuss possible extensions and modifications of the design.
+The ConvCoTM accelerator’s performance is discussed in Section VII, and Section VIII concludes the paper.
 II Related Work
@@
 In [7] a digital 10 nm FinFET CMOS solution is reported.
-Utilizing a BNN algorithm, it achieves a peak energy efficiency of 617 TOPS/W (Tera operations per second per watt).
+Utilizing a BNN algorithm, it achieves a peak energy efficiency of 617 Tera operations per second per watt (TOPS/W).
 It has a power consumption of 5.6 mW and obtains an accuracy of 86% when operating on the CIFAR-10 dataset.
@@
 The main advantage of the ternary approach is a reduction of the required operations per classification with 3.9$\times$ compared to a BNN model.
-The accelerator classifies MNIST images at a rate of 549 FPS, consuming 96 $\mu$W.
-This corresponds to an EPC of 0.18 $\mu$J.
+The accelerator classifies MNIST images at a rate of 549 FPS, consuming 96 $\mu$W, which corresponds to an EPC of 0.18 $\mu$J.
 The test accuracy achieved is 97.1%.
@@
 On MNIST it achieves a test accuracy of 96.92% and 1.38 nJ EPC.
+A time-domain IMC mixed-signal SNN is reported in [26].
+It is implemented in a 65 nm CMOS technology and has an energy efficiency of 701.7 TOPS/W.
+On the CIFAR-10 dataset it achieves a test accuracy of 91.3%.
+In [27] an analog IMC accelerator suited for CNNs as well as for Transformers is presented.
+It achieves 818–4094 TOPS/W, and CIFAR-10 test accuracies of 91.7% and 95.8% for CNN (ResNet20) and Transformer (ViT-S) respectively.
+The technology used is 65 nm CMOS.
 For TM-based HW solutions, the first reported ASIC is described in [11].
@@
 It represents a highly compact solution when it comes to HW resource usage.
-In [12] the first reported ConvCoTM-based HW solution is presented, of which the inference part is the basis for our ASIC as described in the Introduction section.
+In [12] the first reported ConvCoTM-based HW solution is presented, of which the inference part is the basis for our ASIC, as described in the Introduction section.
 It is an FPGA solution that classifies $28\times 28$ images with 10 categories, and includes full on-device training.
@@
 The source code for this design is available at [30].
+The FPGA-based accelerator in [31] is a flexible solution that supports on-device training for the vanilla TM and the CoTM.
+It requires less HW resources and provides greater flexibility compared to [12], at the expense of higher latency.
 In [32, 33, 34] the design and simulation of low-latency asynchronous TM solutions for inference and training are presented.
@@
 III-A General TM Concepts
-The input feature vector to a TM consists of $o$ propositional Boolean variables, $x_{u}\in\{0,1\}^{o}$, where $u\in\{0,\ldots,o-1\}$.
+The input feature vector, $X$, to a TM consists of $o$ Boolean variables, $X=[x_{0},x_{1},\dots,x_{o-1}]$.
 A new input vector named the literals, L, is generated by appending the negated variables to the input, as shown in Eq.
 (1).
-Thus, there are in total $2o$ literals.
+Thus, there are in total $2o$ single-bit literals.
 $L=[l_{0},l_{1},\dots,l_{2o-1}]\\ =[x_{0},\dots,x_{o-1},\neg x_{0},\dots,\neg x_{o-1}]$ (1)
 A TA of two-action type [37] will either include or exclude a literal in a conjunctive clause, depending on its state.
+Fig.
+1 shows the structure of a single TA with $2N$ states.
+In HW, a TA is typically implemented as a binary up/down counter, and the inverted version of its most significant bit (MSB) is used as the TA action signal (active high).
+Figure 1: A Tsetlin Automaton (TA) for two-action environments.
 One team of TAs is applied per clause, and the different TA actions are obtained during the training process, which involves feedback mechanisms to each TA [10].
@@
 where $l_{k}$ is the literal with index k.
-$k$ belongs to $I_{j}\subseteq\{0,\dots,2o-1\}$, where $I_{j}$ is the set of indices of the literals in clause $j$ for which the corresponding TAs take the action include.
+$k$ belongs to the set $I_{j}\subseteq\{0,\dots,2o-1\}$, which is obtained from training, or from a pre-trained model.
+$I_{j}$ consists of the indices of the literals in clause $j$ for which the corresponding TAs take the action include.
 III-B CoTM
@@
 No multiplications are needed to implement Eq.
-(3), only additions, as a clause can output either 0 or 1.
-The predicted class, $\hat{y}$, is determined by the argmax operator, as given by Eq.
-(4).
+(3), as a clause’s output is either 0 or 1.
+The predicted class, $\hat{y}$, is determined by the argmax operator:
 $\hat{y}=\text{argmax}_{i}\{v_{i}\},~{}i\in\{0,\ldots,m-1\}.$ (4)
@@
 In Eq.
-(5), $U$ is the number of bits used for thermometer encoding [35] of the value of a single pixel (one-hot encoding can also be used).
+(5), $N_{F}$ corresponds to $o$, which is the number of Boolean variables which are input to the TM, as described in Subsection III-A.
+$U$ is the number of bits used for thermometer encoding [38] of the value of a single pixel (one-hot encoding can also be used).
 Such encoding is required because a TM operates on Boolean variables, not on binary numbers.
@@
 III-D Datasets and Booleanization
-The widely used MNIST [14] dataset consists of handwritten digits.
-FMNIST [15] is more challenging than MNIST and includes images from a Zalando catalog, such as t-shirts, sandals, and shoes.
-KMNIST [16] consists of cursive Japanese characters, and is also more demanding than MNIST.
+The widely used MNIST dataset consists of handwritten digits.
+FMNIST is more challenging than MNIST and includes images from a Zalando catalog, such as t-shirts, sandals, and shoes.
+KMNIST consists of cursive Japanese characters, and is also more demanding than MNIST.
 Each of these datasets includes images of $28\times 28$ pixels organized in 60k training samples and 10k test samples.
@@
 The convolution window has dimensions $W_{X}=W_{Y}=10$, and a step size of one is used in both directions.
-For each patch, we have 100 feature bits from the convolution window plus 36 bits for the appended patch position information, ref.
+For each patch, we therefore have 100 feature bits from the convolution window plus 36 bits for the appended patch position information, ref.
 Eq.
 (5).
-The number of literals for each patch is therefore 272.
+Thus, there are 272 literals per patch.
 III-E TM and Convolution - Inference
@@
 The clause output is given by Eq.
-(6) which we call the sequential OR-function.
+(6) which we denote as the sequential OR-function.
 Here $c_{j}$ is the $j$’th clause of the ConvCoTM, and $c_{j}^{b}$ is the $b$’th output of this clause obtained during the patch generation:
@@
 The block diagram of the ConvCoTM accelerator is shown in Fig.
-1.
+2.
 The accelerator includes all circuitry required for performing inference, and only minimal interaction with the system processor is required.
@@
 It is used for transferring model data to the accelerator during load model mode, and for transferring images in inference mode.
-When a classification operation is finished, the accelerator outputs an interrupt.
-In addition, it outputs a separate 8-bit signal consisting of the predicted class of the sample together with the true label (if provided).
+When a classification operation is finished, the accelerator outputs an interrupt, and a separate 8-bit signal consisting of the predicted class of the sample together with the true label (if provided).
 IV-B ConvCoTM Model Registers
@@
 The patch generation module consists of a register with 10 rows, each with 28 DFFs.
-At the start of processing a new sample, the first 10 rows of an image are loaded into the register, and the convolution window starts at the leftmost position, which has coordinates ($0,0$).
+At the start of processing a new sample, the first 10 rows of an image are loaded into the register, and the convolution window starts at the leftmost position, which has coordinates ($0,0$), see Fig 3.
+Figure 3: Register structure for window sliding. Each square represents a DFF. The first 10 rows of a booleanized image have been loaded and the window is in the start position (0,0).
 During the patch generation, the window is moved to the right with increments of one.
@@
 There are 19 different positions also for the $y$-coordinate, and during a complete patch generation phase, $19\times 19=361$ patches are generated.
+The $x$- and $y$-positions are thermometer encoded, each with 18 bits as shown in Table I.
+TABLE I: Thermometer position encoding of $10\times 10$ convolution window in a $28\times 28$ image. $x$ at position 0 means leftmost column, while $y$ at position 0 means uppermost row.
+$x$ or $y$ position Thermometer encoded value (18 bits)
+0 000000000000000000
+1 000000000000000001
+$\cdots$ $\cdots$
+17 011111111111111111
+18 111111111111111111
 IV-D Clause Pool
@@
 In this case, the combinational clause logic that generates the signal $c_{j}^{b}$ in Fig.
-2, would have to be replicated for each convolution window.
+4, would have to be replicated for each convolution window.
+The outputs from each combinational clause logic instance would have to be input to the OR-gate labeled A in Fig.
+4.
 IV-E Inference
@@
 (3).
-All ten class sums are generated in parallel, by adders configured in a tree-structure and utilizing a three-stage pipeline.
-Finally, an argmax module [11] selects the class with the highest class sum, as the predicted class, $\hat{y}$.
+All ten class sums are generated in parallel, and each class sum is obtained by adding 128 numbers as shown in Fig.
+5.
+For each adder input, a multiplexer (MUX) is employed, which outputs the clause weight, $w_{i,j}$, if $c_{j}=1$, or zero if $c_{j}=0$.
+The adders are configured in a reduction-tree, with a three-stage pipeline.
+Figure 5: Class sum generation for class $i$. The main input signals are the class weights, $w_{i,j}$, and the Boolean clause values, $c_{j}$, where $j\in\{0,\dots,127\}$.
+Finally, an argmax module selects the class with the highest class sum, as the predicted class $\hat{y}$.
 Fig.
-3 shows a simplified state diagram of the accelerator.
+6 shows the argmax module which is implemented by combinational logic, using a submodule from [11] in a reduction-tree structure.
+In Fig.
+6, this submodule is shown in the upper right part.
+The inputs are two class sums, $v_{0}$ and $v_{1}$, and their corresponding 4-bit class labels, label0 and label1.
+If $v_{1}>v_{0}$, the module feeds forward $v_{1}$ as the $v_{max}$ result, and label1 is output as the maximum argument, $a_{max}$.
+Otherwise, $v_{0}$ and label0 are selected.
+Figure 6: The argmax module, that selects the label corresponding to the largest class sum. Based on a submodule circuit from [11].
+Fig. 7 shows a simplified state diagram of the accelerator, and pseudo-code for the inference procedure is described in Algorithm 1.
+Algorithm 1 ConvCoTM Accelerator Inference
+1:procedure Classify($A$) $\triangleright$ $A$ is a single image example fed to the accelerator by the system processor.
+2:    $\triangleright$ From here the algorithm is performed in its entirety by the accelerator.
+3:    Reset clause output register.$\triangleright$ Resets the DFF in each instance of the clause logic, ref. Fig. 4.
+4:    for $i=0$ to $B-1$ do $\triangleright$ Process $B$ patches, ref. Subsections III-C and IV-C.
+5:       Perform in parallel during a single clock cycle:
+6:       * Generate patch($i$) of $A$
+7:       * Evaluate all clauses, $c_{j}$, $j\in\{0,\dots,n-1\}$, for patch($i$)
+8:       * OR the clause output register $c_{j}$ with the combinational clause output $c_{j}^{b}$
+9:           and update clause output register $\triangleright$ Ref. Eq. (6) and Fig. 4.
+10:    end for
+11:    Calculate class sums: $v_{0},v_{1},\dots,v_{m-1}$triangleright Ref. Eq. (3). A tree structure of adders, pipelined in 3 stages, is used per class sum.
+12:    return argmax{$v_{0},v_{1},\dots,v_{m-1}$} $\triangleright$ Returns predicted class. Ref. Eq. (4).
+13:end procedure
 Figure 7: Simplified accelerator state diagram.
@@
 It includes 99 clock cycles for transferring the 98 image bytes and the label byte, and 372 clock cycles for the patch generation, class summation and class prediction.
-When the accelerator operates in continuous mode, a new sample is loaded into the image buffer while the current sample is being processed, see Subsection IV-C.
+When the accelerator operates in continuous mode, a new sample is loaded into the image buffer while the current sample is being processed, see Subsection IV-C and Fig.
+8.
 This increases the throughput, as image samples can be processed every 372’th clock cycle.
 Any timing overhead in the system processor will add to the total latency.
+Figure 8: Simplified accelerator timing diagram.
 IV-F Clock Domains
@@
 The clock-gating can be enabled/disabled by an external pin.
+V Implementation and Measurement Results
 Figure 9: Photo and layout plot of the ConvCoTM accelerator chip.
-V Implementation and Measurement Results
 The ConvCoTM accelerator was designed in VHDL and implemented in a 65 nm low-leakage CMOS technology from UMC.
+We used the Cadence IC design framework for the chip implementation, and the VHDL simulations were performed with Xcelium.
+Synthesis and Place & Route were performed with Genus and Innovus respectively.
+To reduce leakage current and to optimize the power consumption for low operating frequencies, relaxed time-constraints were applied, and digital cells with high drive strengths were excluded.
+For voltage-drop analysis we applied Voltus, and Conformal was used for logic equivalence checking.
+For timing verification of the routed design, we used mixed VHDL/Verilog gate-level simulations with back-annotated timing information.
+Complete gate-level simulations of the accelerator with model loading and inference of the full MNIST test dataset were performed.
+Design-rule check (DRC) and layer versus schematic (LVS) were performed with Calibre, while GDSII file generation was performed with Virtuoso.
+The VHDL source code for the ASIC implementation (used for initial FPGA verification) is publicly available at a GitHub repository [40].
 A JLCC44 ceramic package housed the chip.
+For the package pins, there were 5 ground connections, 2 supply voltage connections (3.3 V) for the digital inputs/outputs (IOs), 2 supply connections (1.2 V) for the core cells, and 29 IOs.
+Six pins were not connected.
+The number of pins could have been reduced significantly, but as the ASIC was a test chip, we prioritized testability and observability.
 Fig.
@@
 The accelerator’s core area is approximately 2.7 mm2.
-The VHDL source code for the ASIC implementation (used for initial FPGA verification) is available at a GitHub repository [37].
-A printed circuit board (PCB) was developed for testing and characterization of the ASIC.
+For the testing and characterization of the ASIC, a printed circuit board (PCB) was developed.
 The ASIC PCB was connected to a Digilent Zybo Z7-20 development board, equipped with a Xilinx Zynq XC7Z020 FPGA, as shown in Fig.
-5.
+10.
 One of the FPGA’s ARM9 cores was configured as system processor, and the FPGA board provided clocks, model data, image data samples and control signals to the ASIC PCB.
 The processor also read back accelerator status signals, and the predicted class per sample.
-To interface with the FPGA board, the ASIC’s input/output (IO) ports operated at 3.3 V supply voltage.
+To interface with the FPGA board, the ASIC’s IOs operated at 3.3 V supply voltage.
 The accelerator core operated from 0.82 V to 1.2 V.
@@
 These results are independent of the accelerator clock frequency and are exactly in accordance with the performance of the models obtained from the SW simulations.
-TABLE I: The ConvCoTM accelerator ASIC’s main characteristics and performance
+TABLE II: The ConvCoTM accelerator ASIC’s main characteristics and performance
 Parameter / Characteristics Value / Description
@@
 With a supply voltage of 0.82 V, the accelerator core’s power consumption was 0.52 mW and the corresponding EPC was 8.6 nJ.
-With 1.2 V core supply voltage and a clock frequency of 1 MHz, the achieved classification rate was 2.27k FPS at a power consumption of 81 $\mu$W.
-Table I summarizes the characteristics and performance of the accelerator.
-In a real application, the ConvCoTM design would typically be included as a peripheral module to a microcontroller in an SoC.
+Table II summarizes the characteristics and performance of the accelerator.
+In a real application, the ConvCoTM design would typically be included as a peripheral module to a microcontroller in an SoC, and the digital IO interface would have been optimized.
 Therefore, we considered the power consumption of the accelerator core as the most interesting power parameter.
 In inference mode, the digital 3.3 V IOs, which were required for interfacing to the FPGA board, had a power consumption of 0.76 mW, at a clock frequency of 27.8 MHz.
-All power and energy measurement results of the ASIC are average values, and were performed with clock-gating and CSRF enabled (as shown in Fig.
-2).
+All reported power and energy measurement results of the ASIC are average values, and were performed with clock-gating and CSRF enabled (as shown in Fig.
+4).
 When operating at 27.8 MHz, clock-gating reduced the power consumption by approximately 60%, while the CSRF alone provided less than 1% power reduction.
-VI Discussions
-We consider the ConvCoTM accelerator with 128 clauses as a small configuration.
-In light of this, we are satisfied with the obtained test accuracy which is comparable with other low-complexity and ULP solutions.
-Table II shows the performance and properties of state-of-the-art ULP IC implementations for classification of MNIST and CIFAR-10 datasets.
-Inspired by the TinyML [39] benchmarks for embedded systems, we specify the classification rate, the peak power consumption and the EPC.
-These are system parameters which are independent of the accelerator model and architecture [40].
-However, when used for comparing solutions, one must be careful to specify the ML task performed.
-The CIFAR-10 dataset is significantly more challenging than MNIST.
-Therefore, accelerators that operate on CIFAR-10, have lower classification rates and higher EPC compared to accelerators that operate on MNIST.
-TABLE II: Comparison of the ConvCoTM IC accelerator with previous works.
-This work TCAS-I´25 Zhao [20] TCAS-II’23 Yejun [21] JSSC’23       Yang [9] TCAS-I’20 Mauro [6] JSSC’21       Knag [7] TCAS-I’20 Bankman [5]
-Technology 65 nm CMOS 28 nm CMOS 65 nm CMOS 40 nm CMOS 22 nm FD-SOI 10 nm FinFET 28 nm CMOS
-Active area 2.7 mm2 0.261 mm2 0.57 mm2 0.98 mm2 2.3 mm2 0.39 mm2 4.6 mm2
-Algorithm ConvCoTM CNN SNN Ternary CNN BNN BNN BNN
-Design type Digital Analog, time domain Neuromorphic mixed-signal IMC, mixed-signal Digital (SoC) Digital IMC, mixed-signal
-Image dataset MNIST, FMNIST, KMNIST MNIST MNIST MNIST CIFAR-10 CIFAR-10 CIFAR-10
-Measured test accuracy 97.42%, 84.54%, 82.55% 97.9% 95.35% 97.1% 99% of nominal accuracy 86% 86%
-Classifications per second 60.3k a           2.27k b 3508 233k (1.2 V)           40k (0.7 V) 549 15.4 Not stated 237
-Latency e 25.4 $\mu$s a           0.66 ms b Not stated Not stated Not stated Not stated Not stated Not stated
-Inference power 1.15 mW a,c 0.52 mW a,d 81 $\mu$b,c 21 $\mu$b,d 11.6 $\mu$W 9.37 mW (1.2 V) 0.517 mW (0.7 V) 96 $\mu$W 674 $\mu$f 5.6 mW (13 MHz) 607 mW (622 MHz) 0.9 mW
-Energy per classification (EPC) 19.1 nJ a,c           8.6 nJ a,d           35.3 nJ b,c           9.6 nJ b,d 3.32 nJ 40.17 nJ (1.2 V) 12.92 nJ (0.7 V) 0.18 $\mu$J 43.8 $\mu$J Not stated 3.8 $\mu$J
-a 27.8 MHz clock frequency.
-b 1.0 MHz clock frequency.
-c The accelerator core supply voltage (vdd) is 1.20 V.
-d vdd=0.82 V.e For processing of a single image sample, including data transfer.
-f Full SoC.
-The mixed-signal IMC-based solution in [9] operates with a peak power consumption of 96 $\mu$W and consumes 0.18 $\mu$J per classification on MNIST, with a test accuracy of  97.1%.
-In [21] a mixed-signal SNN IC accelerator achieves an EPC of 12.92 nJ with 0.7 V supply voltage.
-It utilizes a time-domain echo state network (ESN) [41] approach and thus requires non-standard interfacing with the surrounding system’s resources.
-The test accuracy achieved on MNIST is 95.35%.
-The 28 nm CMOS CNN-based accelerator in [20] utilizes an IMC architecture with analog and time-domain signal processing.
-It achieves an EPC of 3.32 nJ, which is the lowest reported for MNIST.
-Our ConvCoTM accelerator ASIC achieves an EPC of 8.6 nJ at 0.82 V operating voltage.
-This is the second lowest EPC reported, for a manufactured ASIC that operates on MNIST.
-As the ConvCoTM ASIC is fully digital, the classification rate and thereby also the peak power consumption, can be scaled by adjusting the clock frequency.
-At a clock frequency of 1 MHz, the power consumption is 81 $\mu$W.
-Most ULP accelerators are based on analog techniques [20, 21, 9, 5].
-The test accuracy of such solutions is susceptible to variations in process parameters, supply voltage and operating temperature (PVT).
-All-digital solutions, like our ConvCoTM ASIC and [6, 7, 8], avoid such non-ideal effects.
-In addition, these accelerators are easier to integrate in larger SoCs, and can more easily be scaled and re-targeted to other technologies.
-This applies for CNNs as well as TMs.
-In the ConvCoTM accelerator, the combinational clause logic draws only a small amount of energy compared to the clock tree of the inference-core DFFs.
-This is the main reason why the CSRF technique (see Subsection IV-D) did not provide a greater reduction in EPC.
-Other convolutional TM architectures, that operate at higher clock frequencies, process more literals in parallel, and utilize sparsity in TM models, can potentially benefit more from the CSRF technique.
-VII Scaling
+VI Possible Extensions
+In this section, we describe a technique which limits the number of literals that are input to each clause [42].
+This can enable a significant reduction in model size and chip area, and we provide estimates for a 28 nm CMOS implementation of the ConvCoTM accelerator.
+We also discuss how the accelerator could be modified to support on-device training.
+Finally, we describe concepts and estimates for a scaled-up solution intended for operation on the CIFAR-10 dataset.
+VI-A Estimated Accelerator Performance if Implemented in 28 nm CMOS
+In the current ConvCoTM ASIC accelerator, each clause takes all the literals as input, ref.
+Fig.
+4.
+This makes it necessary to store and use also the exclude TA actions of the model.
+However, for a given clause, the literals that are excluded do not contribute to the the clause value generation and therefore represent an overhead of the model.
+TM models are in general highly sparse.
+For example, in the MNIST model used for the ConvCoTM accelerator, 88% of the TA actions are exclude.
+There is active research on how to utilize the TM model sparsity to achieve more compact HW and SW solutions.
+One option is to use a TM training setting that specifies a maximum number of literals per clause [42], with only negligible loss of accuracy.
+This enables a significant reduction of the model size and the associated chip area for storage and processing of the TA actions.
+As an example, we here consider the current ASIC and assume that each clause only takes 10 literals as input.
+To select a single literal, one would need a 272-to-1 multiplexer, and a 9-bit literal address (multiplexer selection word).
+That is, one would need $10\times 9=90$ bits per clause, for storing which literals are included.
+The principle for the clause logic could be as shown in Fig.
+11.
+In this case, an area reduction of $(272-90)/272\approx 67$% of the TA action part of the model is achieved.
+In the current ASIC, the TA action part of the model, and the associated clause logic, constitute about 70% of the accelerator core’s chip area.
+Using the principle in Fig.
+11, one could achieve a total reduction of the core area of approximately 47%.
+Figure 11: Circuitry for a single clause, where only 10 literals are used as input. $s_{0}$ to $s_{9}$ are multiplexer selection signals given by the model.
+A reduction in chip area and EPC can be achieved by implementing the ConvCoTM accelerator in a more advanced technology.
+We here estimate the chip area, power consumption and EPC if a 28 nm CMOS technology was used.
+With Dennard scaling [43], we roughly estimate a general reduction factor of the active chip area of $(28/65)^{2}$.
+In addition, we assume the number of literals per clause is limited to 10 as described above.
+Thus gives an estimated chip area of 0.27 mm2.
+At an operating frequency of 27.8 MHz (used here for comparison as this was the maximum clock rate tested for the manufactured ConvCoTM ASIC), and 0.7 V supply voltage, we roughly estimate a 50% reduction in power consumption compared to the 65 nm chip operating at 0.82 V.
+This results in an EPC of 4.3 nJ which would be close to what is achieved for the SNN accelerator in [20].
+VI-B Extension of the ASIC Design to Support Training
+Based on the ConvCoTM training concept and circuit modules used in the FPGA accelerator in [12], we here describe how the ASIC could be modified to support on-device training.
+During the patch generation phase, one has to perform random patch selection per clause.
+In [12], this is performed based on the reservoir sampling algorithm [44].
+One would need one or more RAM modules for storing all of the 361 generated patches, each consisting of 136 feature bits.
+In addition, a register of 9 bits is required per clause to store the address of the randomly selected patch.
+For each clause, 272 TAs would be needed, i.e., one TA for each literal.
+We here assume that 8-bit TAs are used, based on the structure shown in Fig.
+1.
+In [12], all TAs belonging to a single clause are updated simultaneously.
+Therefore, the TAs can be implemented by parallel operation of 34 single-port RAM modules, each with a word width of 64 bits, supporting 8 TAs.
+The number of rows in the TA RAMs would be 128, corresponding to the number of clauses in the ConvCoTM configuration used.
+The model registers for the weights and the TA action signals, are already included in the ASIC, but to support training, register update logic would be required.
+For random number generation during the reservoir sampling and during the clause updating phase, linear feedback shift-registers (LFSRs) of minimum 16 bits, are needed [12].
+During the clause updating, one LFSR is used for the random decision of whether a clause’s TAs should receive update, and 272 LFSRs are required for the simultaneous updating of all TAs per clause.
+The additional circuit modules required for implementing the training modules is moderate, and we estimate it to approximately 1 mm2.
+For inference, the power consumption should not increase much, provided clock-gating is applied for the specific training modules, and RAM modules with very low leakage current are used for the TAs.
+Reading from and writing to the RAMs consume energy, but these operations would only be performed during training.
+The accelerator in [12] operated at 50 MHz and achieved a training throughput of approximately 40k samples per second.
+For a modified ASIC that supports training, and operates at 27.8 MHz, we estimate a training throughput of 22.2k samples per second.
+VI-C Scaled-Up Solution
 The ConvCoTM accelerator’s architecture described in this paper, is highly suited for scaling, and can be modified to operate on larger and more complex images.
-There is active research on how to apply the TM algorithm for larger and more complex images.
+There is active research on how to apply the TM algorithm on larger and more complex images.
 The TM version that currently achieves the best test accuracy on CIFAR-10 is TM Composites [17, 18].
@@
 For the TA actions, the model for each TM Specialist therefore needs to include 16 ten-bit literal addresses per clause.
-We assume the same 65 nm CMOS technology is used as for the current ConvCoTM ASIC.
-The chip area for the CIFAR-10 accelerator is estimated by multiplying the chip area of the current ASIC (2.7 mm2) with the ratio, $R$, given by the model size of one TM Specialist (32.5 kilobyte) divided by the model size of the ASIC in this paper (5.6 kilobyte).
+We first assume the same 65 nm CMOS technology is used as for the current ConvCoTM ASIC.
+The chip area for the CIFAR-10 accelerator is estimated by multiplying the chip area of the current ASIC (2.7 mm2) with the ratio, $R$, given by the model size of one TM Specialist (32.5 kilobytes) divided by the model size of the ASIC in this paper (5.6 kilobytes).
 We therefore get $R=32.5/5.6\approx 5.8$.
@@
 With a system clock of 27.8 MHz (as for the current ConvCoTM ASIC), the classification rate would be $808027.8\times 10^{6}/$ 8080 FPS $\approx 3440$ FPS.
-An estimate of the power consumption is found by multiplying the power consumption of the current ASIC’s accelerator core, operating at 27.8 MHz (1.15 mW), with the factor $R$.
+An estimate of the power consumption is found by multiplying the power consumption of the current ASIC’s accelerator core, operating at 27.8 MHz and 0.82 V, with the factor $R$.
 During model loading and booleanization, we assume power consumption similar to that during inference.
-An estimate of the energy per classification is then $R\times 1.15$ mW/(3440 FPS) $\approx 1.9~{}\mu$J/frame.
+An estimate of the EPC is then $R\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.52}$ mW/(3440 FPS) $\approx{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.9}~{}\mu$J/frame.
+For a 28 nm CMOS implementation of this solution, we roughly estimate an accelerator core area of 3.3 mm2, based on Dennard scaling.
+Assuming 0.7 V supply voltage, a coarse estimate of the EPC is 0.45 $\mu$J.
+The operation frequency could be increased for higher throughput, and this could enable further reduction of the EPC.
 TABLE III: Envisaged ConvCoTM inference ASIC Configuration and Performance for operation on the CIFAR-10 Dataset, based on the TM Composites architecture.
@@
 Model size: TA actions 20 1000 clauses, each with 16
-per TM Specialist kilobyte literals. The address
+per TM Specialist kilobytes literals. The address
 length per literal is 10 bits.
-Model size: Weights 12.5k 10 classes, 1000 clauses
-per TM Specialist kilobyte and 10-bit weights.
+Model size: Weights 12.5 10 classes, 1000 clauses
+per TM Specialist kilobytes and 10-bit weights.
 Complete model size 130 For four TM Specialists.
-kilobyte
+kilobytes
 Classification rate 3440 FPS Sequential operation of the
@@
 Accelerator core 17.7 mm2 65 nm CMOS.
-Accelerator core 6.7 mW Inference mode,
-power consumption 27.8 MHz system clock.
-Accelerator core 1.9 $\mu$J Inference mode,
-energy per 27.8 MHz system clock.
-classification.
+3.3 mm2 28 nm CMOS.
+Accelerator core 3.0 mW 27.8 MHz, 0.82 V, 65 nm.
+power consumption 1.5 mW 27.8 MHz, 0.7 V, 28 nm.
+Accelerator core 0.9 $\mu$J 27.8 MHz, 0.82 V, 65 nm.
+energy per 0.45 $\mu$J 27.8 MHz, 0.7 V, 28 nm.
+classification (EPC).
+VII Discussions
+TABLE IV: Comparison of the ConvCoTM IC accelerator with previous works.
+This work Envisaged design. This work scaled to 28 nm. TCAS-I´25 Zhao [20] TCAS-II’23 Yejun [21] JSSC’23       Yang [9]
+Technology 65 nm CMOS 28 nm CMOS 28 nm CMOS 65 nm CMOS 40 nm CMOS
+Active area 2.7 mm2 0.27 mm2 0.261 mm2 0.57 mm2 0.98 mm2
+Algorithm ConvCoTM ConvCoTM CNN SNN Ternary CNN
+Design type Digital Digital Analog, time domain Neuromorphic mixed-signal IMC, mixed-signal
+Image dataset MNIST, FMNIST, KMNIST MNIST, FMNIST, KMNIST MNIST MNIST MNIST
+Measured test accuracy 97.42%, 84.54%, 82.55% 97.42%, 84.54%, 82.55% 97.9% 95.35% 97.1%
+Classifications per second 60.3 k a           2.27 k b 60.3 k a 3508 233 k (1.2 V)           40 k (0.7 V) 549
+Latency e 25.4 $\mu$s a           0.66 ms b 25.4 $\mu$s a Not stated Not stated Not stated
+Inference power 1.15 mW a,c 0.52 mW a,d 81 $\mu$b,c 21 $\mu$b,d 0.26 mW a,f 11.6 $\mu$W 9.37 mW (1.2 V) 0.517 mW (0.7 V) 96 $\mu$W
+Energy per classification (EPC) 19.1 nJ a,c           8.6 nJ a,d           35.3 nJ b,c           9.6 nJ b,d 4.3 nJ a,f 3.32 nJ 40.17 nJ (1.2 V) 12.92 nJ (0.7 V) 0.18 $\mu$J
+a 27.8 MHz clock frequency.
+b 1.0 MHz clock frequency.
+c The accelerator core supply voltage (vdd) is 1.20 V.
+d vdd=0.82 V.
+e For processing of a single image sample, including data transfer.
+f vdd=0.7 V.
+TABLE V: Comparison of the envisaged scaled-up ConvCoTM IC accelerator with previous works that operate on CIFAR-10.
+Envisaged design from Subsection VI-C TCAS-I’20 Mauro [6] JSSC’21       Knag [7] TCAS-I’20 Bankman [5] TCAS-I’25 Park [26] JSSC’25 Yoshioka [27]
+Technology 65 nm CMOS 28 nm CMOS 22 nm FD-SOI 10 nm FinFET 28 nm CMOS 65 nm CMOS 65 nm CMOS
+Active area 17.7 mm2 (65 nm) 3.3 mm2 (28 nm) 2.3 mm2 0.39 mm2 4.6 mm2 0.17 mm2 0.48 mm2
+Algorithm ConvCoTM BNN (SW defined) BNN BNN SNN (spiking VGG-16) CNN / Transformer
+Design type Digital Digital (SoC) Digital IMC, mixed-signal Analog, with time-domain IMC Analog, IMC
+Image dataset CIFAR-10 CIFAR-10 CIFAR-10 CIFAR-10 CIFAR-10 CIFAR-10
+Test accuracy 79% 99% of nominal accuracy 86% 86% 91.13% 91.7% j           95.8% k
+Classifications per second 3440 a 15.4 Not stated 237 Not stated Not stated
+Latency b 0.3 ms a Not stated Not stated Not stated Not stated Not stated
+Inference power 3.0 mW a,c (65 nm) 1.5 mW a,d (28 nm) 674 $\mu$e 5.6 mW l 607 mW m 0.9 mW 0.55 mW i Not stated
+TOPS/W Not stated f 13 e,n 617 l         269 m 532 701.7 g         488.7 h 4094 j         818 k
+Energy per classification (EPC) 0.9 $\mu$a,c (65 nm) 0.45 $\mu$a,d (28 nm) 43.8 $\mu$J Not stated 3.8 $\mu$J Not stated Not stated
+a 27.8 MHz clock frequency.
+b For processing of a single image sample, including data transfer.
+c vdd=0.82 V.
+d vdd=0.7 V.
+e For the complete SoC.
+f Not stated due to the very different architecture compared to DNNs and SNNs.
+g Macro.
+h System.
+i 50 MHz, 0.75 V.
+j CNN.
+k Transformer.
+l 13 MHz, 0.37 V.
+m 622 MHz, 0.75 V.
+n 0.5 V.
+We consider the ConvCoTM accelerator with 128 clauses as a small configuration.
+In light of this, we are satisfied with the obtained test accuracy which is comparable with other low-complexity and ULP solutions.
+Table IV shows the performance and properties of state-of-the-art ULP IC implementations for classification of MNIST.
+There, we have also included the performance of the envisaged 28 nm CMOS implementation of the ConvCoTM accelerator, described in Subsection VI-A.
+In Table V the envisaged accelerator design from Subsection VI-C is compared with other ULP accelerators that operate on the CIFAR-10 dataset.
+Inspired by the TinyML [46] benchmarks for embedded systems, we specify the classification rate, the peak power consumption and the EPC.
+These are system parameters which are independent of the accelerator model and architecture [47].
+However, when used for comparing solutions, one must be careful to specify the ML task performed.
+The CIFAR-10 dataset is significantly more challenging than MNIST.
+Therefore, accelerators that operate on CIFAR-10, have lower classification rates and higher EPC compared to accelerators that operate on MNIST.
+The mixed-signal IMC-based solution in [9] operates with a peak power consumption of 96 $\mu$W and consumes 0.18 $\mu$J per classification on MNIST, with a test accuracy of  97.1%.
+In [21] a mixed-signal SNN IC accelerator achieves an EPC of 12.92 nJ with 0.7 V supply voltage.
+It utilizes a time-domain echo state network (ESN) [48] approach and thus requires non-standard interfacing with the surrounding system’s resources.
+The test accuracy achieved on MNIST is 95.35%.
+The 28 nm CMOS CNN-based accelerator in [20] utilizes an IMC architecture with analog and time-domain signal processing.
+It achieves an EPC of 3.32 nJ, which is the lowest reported for MNIST.
+Our ConvCoTM accelerator ASIC achieves an EPC of 8.6 nJ at 0.82 V operating voltage.
+This is the second lowest EPC reported, for a manufactured ASIC that operates on MNIST.
+As the ConvCoTM ASIC is fully digital, the classification rate and thereby also the peak power consumption, can be scaled by adjusting the clock frequency.
+At a clock frequency of 1 MHz and a supply voltage of 0.82 V, the achieved classification rate was 2.27 k FPS at a power consumption of 21 $\mu$W.
+Most ULP accelerators are based on analog techniques [20, 21, 9, 5].
+The test accuracy of such solutions is susceptible to variations in process parameters, supply voltage and operating temperature (PVT).
+All-digital solutions, like our ConvCoTM ASIC and [6, 7, 8], avoid such non-ideal effects.
+In addition, these accelerators are easier to integrate in larger SoCs, and can more easily be scaled and re-targeted to other technologies.
+This applies for CNNs as well as TMs.
+In the ConvCoTM accelerator, the combinational clause logic draws only a small amount of energy compared to the clock tree of the inference-core DFFs.
+This is the main reason why the CSRF technique (see Subsection IV-D) did not provide a greater reduction in EPC.
+Other convolutional TM architectures, that operate at higher clock frequencies, process more literals in parallel, and utilize sparsity in TM models, can potentially benefit more from the CSRF technique.
+The estimated EPC of the envisaged design described in Subsection VI-C, for operation on CIFAR-10 images, is very low.
+However, the test accuracy achieved with the TM algorithm for multi-channel images, is still not at the same level as for CNNs, BNNs and SNNs.
+Further research is required to obtain improvement on this performance parameter.
+The accelerators in [7, 26, 27] achieve very high performance in terms of TOPS/W.
+However, as the classification rates are not specified, one cannot easily calculate the EPC numbers for these solutions.
+An overview of select TM-based HW solutions, including this work, is shown in Table VI.
+The solutions span from all-digital to mixed-signal IMC.
+There are two silicon-proven ASICs and many FPGA solutions.
+Several TM concepts have been evaluated based on simulations.
+TABLE VI: Overview of TM-based HW solutions.
+This work Phil. Trans. A’20 Wheeldon [11] TCAS-I’25 Mao [31] TCAS-I’25 Tunheim [12] ISTM’23 Sahu [29] MICPRO’23 Tunheim [28] ISLPED’23 Ghazal [35] Phil. Trans. A’25 Ghazal [36]
+Technology and design status ASIC, 65 nm CMOS, silicon proven ASIC, 65 nm CMOS, silicon-proven FPGA FPGA FPGA FPGA ASIC, simulation ASIC, simulation
+Algorithm ConvCoTM Vanilla TM Vanilla TM, CoTM ConvCoTM Vanilla TM CTM Vanilla TM CoTM
+Operation Inference Inference Training and inference Training and inference Inference Training and Inference Inference Inference
+Design type Digital Digital Digital Digital Digital Digital IMC (ReRAM), mixed-signal IMC (Y-flash), mixed-signal
+Dataset MNIST, FMNIST, KMNIST Binary IRIS MNIST, FMNIST, KMNIST MNIST, FMNIST, KMNIST MNIST 2D Noisy       XOR c MNIST, FMNIST, KMNIST, KWS-6 d MNIST
+Test accuracy 97.42%, 84.54%, 82.55% 97.0% 97.74%, 86.38%, 83.11% 97.6%,   84.1%,   82.8% 97.71% 99.9% 96.48%, 87.67%, 88.6%, 87.1% 96.3%
+Classifications per second 60.3 k a Not stated 22.4 k 134 k Not stated f 4.4 M Not stated Not stated
+Inference power 0.52 mW a,b Not stated 1.65 W 1.8 W Not stated 2.529 W Not stated Not stated
+Energy per classification 8.6 nJ a,b Not stated e 73.6 $\mu$J 13.3 $\mu$J Not stated 0.6 $\mu$J 13.9 nJ, 23.66 nJ, 26.47 nJ, 5.91 nJ Not stated g
+a 27.8 MHz clock frequency.
+b vdd=0.82 V c $4\times 4$ booleanized images.
+d Keyword spotting dataset with six words.
+e 62.7 TOP/J.
+f The number of clock cycles required per classification is stated, for different number of clauses and with/without literal budgeting.
+g 24.56 TOPS/W.
 VIII Conclusion